### v1

# Random covering by rectangles on self-similar carpets

Abstract.
In thus article, given a base-$b$ self-similar set $K$, we study the random covering of $K$ by horizontal or vertical rectangles, with respect to the Alfhors-regular measure on $K$, and the rectangular shrinking target problem on $K$.
1. Introduction
The metric approximation theory aims at estimating the dimension of sets of elements which are approximable at a certain rate by a sequence of particular points of interest. More precisely, given $d\in\mathbb{N}$, $(x_{n})_{n\in\mathbb{N}}\in\Big(\mathbb{R}^{d}\Big)^{\mathbb{N}}$ and $\psi:\mathbb{N}\to\mathbb{R}_{+}$ a mapping satisfying $\lim_{n\to+\infty}\psi(n)=0$, the set $E_{\psi}$ of elements approximable at rate $\psi$ by the sequence $(x_{n})_{n\in\mathbb{N}}$ is defined as
[…]
where $i.o.$ (`` infinitely often ") means that the inequality holds for infinitely many $n$. Such problem were originally born in Diophantine approximation, where one studies sets of real numbers or vectors approximable by rational numbers or vectors at a given speed rate. Such questions are also natural in multifractal analysis, as for many mappings, the regularity at a given point depends of its rate of approximation by specific dyadic numbers or rational numbers (see [19] for instance) and in dynamical and random approximation, as, for instance, given an ergodic system $(T,\mu),$ the local dimension at a given point $x$ depends on its approximation rate by typical $\mu$ orbits, provided that $\mu$-mixes sufficiently fast (see [16, 10]). In particular, the theory of random approximation as raised many interests these last 30 years and has known recently many developments. Given $\mu\in\mathcal{M}(\mathbb{R}^{d})$ a probability measure and $(X_{n})_{n\in\mathbb{N}}$ an i.i.d. sequence of random variable of law $\mu$ it was established in [20] that, denoting $\overline{\dim}_{H}\mu$ the upper Hausdorff dimension of $\mu$ (see Definition 2.2), for any $\delta\geq\frac{1}{\overline{\dim}_{H}\mu},$ almost surely, one has
[…]
This result was later on generalized in [21], showing that no simple formula (depending on geometric quantity related to $\mu$ one usually considers) holds in general in the case $\delta<\frac{1}{\overline{\dim}_{H}\mu}$, solving a conjecture of Eckström and Persson stated in [11]. The question of random approximation by other shapes than ball has also been considered. An important result regarding this topic was established in [14] in the case of measure which are not purely singular. In particular, the authors proved that, given $(X_{n})_{n\in\mathbb{N}}$ i.i.d. of law the Lebesgue measure $\mathcal{L}^{d}$ on $\mathbb{T}^{d}$ and $(O_{n})_{n\in\mathbb{N}}$ a sequence of open sets satisfying $|O_{n}|\to 0,$ one has almost surely,
(2) $\dim_{H}\limsup_{n\to+\infty}\Big(X_{n}+O_{n}\Big)=\inf\left\{t:\sum_{n\geq 1}\mathcal{H}^{t}_{\infty}(O_{n})<+\infty\right\},$
where $\mathcal{H}^{t}_{\infty}(O_{n})$ denotes the Hausdorff content of dimension $s$ of $O_{n}$ (Definition 5).
In the case where the measure $\mu$ is singular, very little is known about the approximation by specific sequences of open sets. Of course, there is no reason, a priori, for a tractable formula to hold as in (2), if the sets $(O_{n})_{n\in\mathbb{N}}$ do not enjoy special properties with respect to $\mu.$ In the present article, we study the random approximation by horizontal (or vertical) rectangles on a self-similar carpet. Let $b\in\mathbb{N}$ be an integer and let $K\subset[0,1]^{2}$ be a base $b$-missing digit set (see Definition 3.1), $\mu_{0}$ the Alfhors-regular measure on $K$ and $(X_{n})_{n\in\mathbb{N}}$ an i.i.d. sequence of law $\mu_{0}.$ Let us fix also $\frac{1}{\dim_{H}K}\leq\tau_{1}\leq\tau_{2}$, $\tau=\frac{\tau_{2}}{\tau_{1}}$ and define
$W_{\tau_{1},\tau_{2}}=\limsup_{n\to+\infty}\Big(X_{n}+(-\frac{1}{n^{\tau_{1}}},\frac{1}{n^{\tau_{1}}})\times(-\frac{1}{n^{\tau_{2}}},\frac{1}{n^{\tau_{2}}})\Big)$
and for $\alpha\geq 0,$ write
$v_{\tau}(\alpha)=\dim_{H}K+(\tau-2)\alpha-(\tau-1)D_{\pi_{2}\mu_{0}}(\alpha),$
where $\pi_{2}\mu_{0}$ denotes the projection of $\mu_{0}$ along the $y$-axis and $D_{\pi_{2}\mu_{0}}$ its multifractal spectrum (Definition 2.4). Let $\beta_{\tau_{1},\tau_{2}}$ be the smallest solution (when well defined) of $v_{\tau}(\beta)=\frac{1}{\tau}.$
Then, there exists $\kappa_{2}\geq 0$ such that, almost surely
$\dim_{H}W_{\tau_{1},\tau_{2}}=\begin{cases}\frac{1}{\tau_{1}}\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \text{ if }\ \frac{1}{\tau_{1}}\leq\dim_{H}\mu_{0}-\dim_{H}\pi_{2}\mu_{0},\\ \frac{1}{\tau_{1}}-(\tau-1)(\beta_{\tau_{1},\tau_{2}}-D_{\pi_{2}\mu_{0}}(\beta_{\tau_{1},\tau_{2}}))\text{ if }\ \dim_{H}\mu_{0}-\dim_{H}\pi_{2}\mu_{0}\leq\frac{1}{\tau_{1}}\leq v_{\tau}(\kappa_{2}),\\ \frac{1+(\tau_{2}-\tau_{1})(s_{0}-2\kappa_{2}+D_{\pi_{2}\mu_{0}}(\kappa_{2})))}{\tau_{2}}\ \ \ \ \ \ \ \ \ \ \ \text{ if }\ \frac{1}{\tau_{1}}\geq v_{\tau}(\kappa_{2}),\end{cases}$
For a more precise statement, we refer to Theorem 3.1 below.
An other very natural approximation problem of dynamical nature is the shrinking target problem. It was originally defined in [17] when the `` targets " are balls. Given a measurable mapping $T:\mathbb{R}^{d}\to\mathbb{R}^{d}$, $x\in\mathbb{R}^{d}$ and $\psi:\mathbb{N}\to\mathbb{R}_{+}$ it consists in studying
$\dim_{H}\left\{y\in\mathbb{R}^{d}:\ T^{n}(y)\in B(x,\psi(n))\text{ i.o. }\right\}:=E_{\psi}(x).$
When $K$ is a two-dimensional base $b$-missing digit set, associated with the IFS $S=\left\{f_{1},...,f_{m}\right\}$ and $T:K\to K$ is defined by $T(y)=by$, on can rewrite
$E_{\psi}(x)=\limsup_{n\geq 1,(i_{1},...,i_{n})\in\left\{1,...,m\right\}^{n}}B\Big(f_{i_{1}}\circ...\circ f_{i_{n}}(x),\psi(n)b^{-n}\Big)$
and it was established by Beresnevitch and Velani that, for every $x\in K,$ one has
$\dim_{H}E_{\psi}(x)=\frac{\dim_{H}K}{1+\liminf_{n\to+\infty}\frac{\log\psi(n)}{-n\log b}}.$
The shrinking target problem on fractals has known many developments since, one refers for instance to [3, 1, 7, 4] In the present article, we study the shrinking target when ``the targets " are taken to be rectangles rather than balls. Unlike the case of balls, our result will depend, in general, on the choice of $x$, the center of our targets. More precisely, let $\nu$ be a $\times b$-ergodic measure supported on $K$, let $1\leq\tau_{1}\leq\tau_{2}$ be two real numbers and write
$V_{\tau_{1},\tau_{2}}=\limsup_{n\geq 1,(i_{1},...,i_{n})\in\left\{1,...,m\right\}^{n}}\Big(f_{i_{1}}\circ...\circ f_{i_{n}}(x)+(-b^{-\tau_{1}n},b^{-\tau_{1}n})\times(-b^{-\tau_{2}n},b^{-\tau_{2}n})\Big).$
Notice that
$V_{\tau_{1},\tau_{2}}=\left\{y\in K:\ T^{n}(y)\in\Big(x+(-b^{-(\tau_{1}-1)n},b^{-(\tau_{1}-1)n})\times(-b^{-(\tau_{2}-1)n},b^{-(\tau_{2}-1)n})\Big)\right\}.$
We prove the following: for $\nu$-almost every $x$, one has
$\dim_{H}V_{\tau_{1},\tau_{2}}(x)=\min\left\{\frac{\dim_{H}K}{\tau_{1}},\frac{\dim_{H}K+(\tau_{2}-\tau_{1})(\dim_{H}K-\alpha_{\nu})}{\tau_{2}}\right\},$
where $\alpha_{\nu}$ is the almost sure local dimension of $\pi_{2}(x)$ with respect to $\pi_{2}\mu_{0}$ (see Proposition 3.3). We refer to Theorem 3.4 for a more general statement and Corollary 3.5 for a formula for general approximation function along the $x$ and $y$-axis. Notice that set of possible dimensions (which are all attained) for $V_{\tau_{1},\tau_{2}}$ when $\nu$ varies in the set of ergodic measures is
$\left\{\min\left\{\frac{\dim_{H}K}{\tau_{1}},\frac{\dim_{H}K+(\tau_{2}-\tau_{1})(\dim_{H}K-\alpha)}{\tau_{2}}\right\},\ \alpha\in\mbox{ Spectr}(\pi_{2}\mu_{0})\right\},$
where  $\mbox{ Spectr}(\pi_{2}\mu_{0})$ denotes the set of possible local dimensions of the measure $\pi_{2}\mu_{0}.$ Finally, we mention that this result regarding the rectangular shrinking targets problems was also established, independently, by Allen, Jordan and Ward in [2] (see Remark 3.6 for more details).
In Section 2, we recall the basis of geometric measure theory, theory of self-similar fractals and multifractal analysis. Our main results regarding the random covering by rectangles and the rectangular shrinking targets problem are stated in Section 3 and the three last sections are dedicated to the proof of these theorems.
2. Preliminaries and notations
Let us start with some notations
Let $d$ $\in\mathbb{N}$. For $x\in\mathbb{R}^{d}$, $r>0$, $B(x,r)$ stands for the closed ball of ($\mathbb{R}^{d}$$\parallel\ \ \parallel_{\infty}$) of center $x$ and radius $r$. Given a ball $B$, $|B|$ stands for the diameter of $B$. For $t\geq 0$, $\delta\in\mathbb{R}$ and $B=B(x,r)$, tBtB stands for $B(x,tr)$, i.e. the ball with same center as $B$ and radius multiplied by $t$, and the $\delta$-contracted ball $B^{\delta}$ is defined by $B^{\delta}=B(x,r^{\delta})$.
[…]
Given a topological space $X$, the Borel $\sigma$-algebra of $X$ is denoted $\mathcal{B}(X)$ and the space of probability measure on $\mathcal{B}(X)$ is denoted $\mathcal{M}(X).$
Given a metric space $X$ and $r>0.$ A $r$-packing of $X$ will consists of a set of open balls $\mathcal{T}$ such that for every $B\in\mathcal{T},$ $|B|=r$ and for every $L\neq B\in\mathcal{T},$ $L\cap B=\emptyset.$
The $d$-dimensional Lebesgue measure on $(\mathbb{R}^{d},\mathcal{B}(\mathbb{R}^{d}))$ is denoted by $\mathcal{L}^{d}$.
For $\mu\in\mathcal{M}(\mathbb{R}^{d})$, $\mathrm{supp}(\mu)=\left\{x\in\mathbb{R}^{d}:\ \forall r>0,\ \mu(B(x,r))>0\right\}$ is the topological support of $\mu$.
Given $X,Y$ two spaces endowed with $\sigma$-algebras and $\mu\in\mathcal{M}(X)$ a measure, for any measurable $f:X\to Y$, one will denote $f\mu\in\mathcal{M}(Y)$ the measure $\mu\circ f^{-1}(\cdot).$
Given $E\subset\mathbb{R}^{d}$, $\dim_{H}(E)$ and $\dim_{P}(E)$ denote respectively the Hausdorff and the packing dimension of $E$.
Given a set $S$, $\chi_{S}$ denotes the indicator function of $S$, i.e., $\chi_{S}(x)=1$ if $x\in S$ and $\chi_{S}(x)=0$ otherwise.
Given $b,n\in\mathbb{N},$ $\mathcal{D}_{b,n}$ or simply $\mathcal{D}_{n}$ when there is no ambiguity on $b$, denotes the set of $b$-adic cubes of generation $n$ and $\mathcal{D}$ the set of all $b$-adic cubes, i.e.
$\mathcal{D}_{n}=\left\{b^{-n}(k_{1},...,k_{d})+b^{-n}[0,1)^{d},\ (k_{1},...,k_{d})\in\mathbb{Z}^{d}\right\}\text{ and }\mathcal{D}=\bigcup_{n\geq 0}\mathcal{D}_{n}.$
In addition, given $x\in\mathbb{R}^{d},$ $D_{b,n}(x)$ or $D_{n}(x)$ will denote the $b$-adic cube of generation $n$ containing $x$.
2.1. Recall on geometric measure theory
Definition 2.1.
Let $\zeta:\mathbb{R}^{+}\mapsto\mathbb{R}^{+}$. Suppose that $\zeta$ is increasing in a neighborhood of 0 and $\zeta(0)=0$. The Hausdorff outer measure at scale $t\in(0,+\infty]$ associated with the gauge $\zeta$ of a set $E$ is defined by
(3) $\mathcal{H}^{\zeta}_{t}(E)=\inf\left\{\sum_{n\in\mathbb{N}}\zeta(|B_{n}|):\,|B_{n}|\leq t,\ B_{n}\text{ closed ball and }E\subset\bigcup_{n\in\mathbb{N}}B_{n}\right\}.$
The Hausdorff measure associated with $\zeta$ of a set $E$ is defined by
(4) $\mathcal{H}^{\zeta}(E)=\lim_{t\to 0^{+}}\mathcal{H}^{\zeta}_{t}(E).$
For $t\in(0,+\infty]$, $s\geq 0$ and $\zeta:x\mapsto x^{s}$, one simply uses the usual notation $\mathcal{H}^{\zeta}_{t}(E)=\mathcal{H}^{s}_{t}(E)$ and $\mathcal{H}^{\zeta}(E)=\mathcal{H}^{s}(E)$, and these measures are called $s$-dimensional Hausdorff outer measure at scale $t\in(0,+\infty]$ and $s$-dimensional Hausdorff measure respectively. Thus,
(5) $\mathcal{H}^{s}_{t}(E)=\inf\left\{\sum_{n\in\mathbb{N}}|B_{n}|^{s}:\,|B_{n}|\leq t,\ B_{n}\text{ closed ball and }E\subset\bigcup_{n\in\mathbb{N}}B_{n}\right\}.$
The quantity $\mathcal{H}^{s}_{\infty}(E)$ (obtained for $t=+\infty$) is called the $s$-dimensional Hausdorff content of the set $E$.
Definition 2.2.
Let $\mu\in\mathcal{M}(\mathbb{R}^{d})$. For $x\in\mathrm{supp}(\mu)$, the lower and upper local dimensions of $\mu$ at $x$ are defined as
$\underline{\dim}_{{\rm loc}}(\mu,x)=\liminf_{r\rightarrow 0^{+}}\frac{\log(\mu(B(x,r)))}{\log(r)}\mbox{ and }\ \ \overline{\dim}_{{\rm loc}}(\mu,x)=\limsup_{r\rightarrow 0^{+}}\frac{\log(\mu(B(x,r)))}{\log(r)}.$
Then, the lower and upper Hausdorff dimensions of $\mu$ are defined by
(6) $\underline{\dim}_{H}(\mu)={\mathrm{ess\,inf}}_{\mu}(\underline{\dim}_{{\rm loc}}(\mu,x))\ \ \mbox{ and }\ \ \overline{\dim}_{P}(\mu)={\mathrm{ess\,sup}}_{\mu}(\overline{\dim}_{{\rm loc}}(\mu,x))$
respectively.
It is known (for more details see [13]) that
$\begin{split}\underline{\dim}_{H}(\mu)&=\inf\{\dim_{H}(E):\,E\in\mathcal{B}(\mathbb{R}^{d}),\,\mu(E)>0\}\\ \overline{\dim}_{P}(\mu)&=\inf\{\dim_{P}(E):\,E\in\mathcal{B}(\mathbb{R}^{d}),\,\mu(E)=1\}.\end{split}$
When $\underline{\dim}_{H}(\mu)=\overline{\dim}_{P}(\mu)$, this common value is simply denoted by $\dim(\mu)$ and $\mu$ is said to be exact-dimensional.
Moreover, a measure $\mu\in\mathcal{M}(\mathbb{R}^{d})$ is called Alfhors-regular if there exists $0\leq\alpha\leq d$ and $C>0$ such that for every $x\in\mathrm{supp}(\mu),$ for every $0<r\leq 1,$ one has
$C^{-1}r^{\alpha}\leq\mu\Big(B(x,r)\Big)\leq Cr^{\alpha}.$
It is direct to check that such a measure is $\alpha$-exact dimensional.
2.2. Self-similar measures and multifractal analysis
Let us start by recalling the definition of a self-similar measure.
Definition 2.3.
A self-similar IFS is a family $S=\left\{f_{i}\right\}_{1\leq i\leq m}$ of $m\geq 2$ contracting similarities of $\mathbb{R}^{d}$.
Let $(p_{i})_{i=1,...,m}\in(0,1)^{m}$ be a positive probability vector, i.e. $p_{1}+\cdots+p_{m}=1$.
The self-similar measure $\mu$ associated with $\left\{f_{i}\right\}_{1\leq i\leq m}$ and $(p_{i})_{1\leq i\leq m}$ is the unique probability measure such that
(7) $\mu=\sum_{i=1}^{m}p_{i}\mu\circ f_{i}^{-1}.$
The topological support of $\mu$ is the attractor of $S$, that is the unique non-empty compact set $K\subset X$ such that $K=\bigcup_{i=1}^{m}f_{i}(K)$.
The existence and uniqueness of $K$ and $\mu$ are standard results [18]. Recall that due to a result by Feng and Hu [15], any self-similar measure is exact dimensional.
2.3. Multifractal analysis of self-similar measure satisfying OSC
Let us start by defining the multifractal spectrum of a measure.
Definition 2.4.
Let $\mu\in\mathcal{M}(\mathbb{R}^{d})$ be a measure and $h\geq 0.$ Set
$E_{h}=\left\{x\in\mathrm{supp}(\mu)\ :\ \lim_{r\to 0^{+}}\frac{\log\mu(B(x,r))}{\log r}=h.\right\}.$
The multifractal spectrum of $\mu$ is the mapping $D_{\mu}$, defined for every $h\geq 0$ by
$D_{\mu}(h)=\dim_{H}E_{h}.$
Moreover, we call
$\mbox{Spectr}(\mu)=\overline{\left\{\alpha:\ D_{\mu}(\alpha)>0\right\}}.$
Let us also recall that a self-similar IFS $S=\left\{f_{1},...,f_{m}\right\}$ is said to satisfy the open set condition if there exists a non empty open set $O$ such that
(8) $\forall 1\leq i\neq j\leq m,\ f_{i}(O)\cap f_{j}(O)=\emptyset.$
Given a self-similar IFS $S$ satisfying the open set condition, we will also say, by extension, that a self-similar measure $\mu$ associated with $S$ satisfies the open set condition.
The multifractal spectrum of any such measure is well understood. Fix $S\left\{f_{1},..,f_{m}\right\}$ a self-similar IFS and $(p_{1},...,p_{m})\in(0,1)^{m}$ a probability vector. For $1\leq i\leq m,$ let $0<c_{i}<1$ be the contraction ratio of $f_{i}$ and, given $q\in\mathbb{R},$ set
$p_{i,q}=c_{i}^{-T(q)}p_{i}^{q},$
where $T(q)$ is such that
$\sum_{1\leq i\leq m}p_{i,q}=1.$
Call also $\mu_{q}$ the self-similar measure associated with $(p_{i,q})_{1\leq i\leq m}$ and
(9) $\begin{cases}\theta_{q}=\frac{\sum_{1\leq i\leq m}p_{i,q}\log p_{i,q}}{\sum_{1\leq i\leq m}p_{i,q}\log c_{i}}\\ \kappa_{q}=\frac{\sum_{1\leq i\leq m}p_{i,q}\log p_{i}}{\sum_{1\leq i\leq m}p_{i,q}\log c_{i}}.\end{cases}$
We recall some of the properties of this spectrum.
Proposition 2.1 ([12], pages 286-295).
Let $\mu\in\mathcal{M}(\mathbb{R}^{d})$ be a self-similar measure satisfying the open set condition. Then:
  • for any $\alpha\geq 0,$ $D_{\mu}(\alpha)\leq\alpha,$
  • the mapping $\alpha\mapsto D_{\mu}(\alpha)$ satisfies is concave and reaches its maximum on $\alpha=\dim_{H}\mu,$
  • $\mbox{Spectr}(\mu)$ is a compact interval. Moreover $\mbox{Spectr}(\mu)=\left\{s\right\}\Leftrightarrow$ $\mu$ is Alfhors-regular,
  • if $\mu$ is not Alfhors-regular, then $D_{\mu}$ is $\mathcal{C}^{\infty}$ on $\overset{\circ}{\mbox{Spectr}(\mu)}=(\alpha_{\min},\alpha_{\max}).$ Moreover $D^{\prime}_{\mu}$ is non increasing on $(\alpha_{\min},\alpha_{\max})$ and
    $\begin{cases}\lim_{\alpha\to\alpha_{\min}}D^{\prime}_{\mu}(\alpha)=+\infty\\ \lim_{\alpha\to\alpha_{\max}}D^{\prime}_{\mu}(\alpha)=-\infty.\end{cases}$
    In addition, for any $q\in\mathbb{R},$ $D_{\mu}(\kappa_{q})=\theta_{q}$, $D^{\prime}_{\mu}(\kappa_{q})=q$ and $\mu_{q}(E_{\kappa_{q}})=1.$
In addition of these properties, when $\mu$ is a self-similar measure satisfying the open set condition, it is also known that the multifractal spectrum and the so-called coarce multifractal spectrum coincides. More precisely, we have the following large deviation estimates.
Theorem 2.2.
Let $\mu\in\mathcal{M}(\mathbb{R}^{d})$ be a self-similar measure satisfying the open set condition. Write $\mbox{Spectr}(\mu)=[\alpha_{\min},\alpha_{\max}].$ Let $\alpha_{\min}\leq\alpha\leq\alpha_{\max}$ be a real number and, given $r>0$ let us write
$\mathcal{P}_{\alpha}(r,\varepsilon)=\sup\#\left\{\mathcal{T}\right\},$
where $\mathcal{T}$ is a $r$-packing of $\mathrm{supp}(\mu)$ with, for every $B\in\mathcal{T}$,
$|B|^{\alpha+\varepsilon}\leq\mu(B)\leq|B|^{\alpha-\varepsilon}.$
There exists $r_{\alpha}>0$ such that for every $r\leq r_{\alpha},$ one has
$r^{-D_{\mu}(\alpha)+\varepsilon}\leq\mathcal{P}_{\alpha}(r,\varepsilon)\leq r^{-D_{\mu}(\alpha)-\varepsilon}.$
3. Main results
3.1. Random covering of self-similar carpet by rectangles
Let us start by defining base-$b$ missing digit IFS’s.
Definition 3.1.
Let $b\in\mathbb{N}$ be an integer and $\mathcal{A}\subset\left\{0,...,b-1\right\}^{2}.$ Given $(i,j)\in\mathcal{A},$ let $g_{(i,j)}$ be the canonical contraction from $0,1]^{2}$ to $(\frac{i}{b},\frac{j}{b})+[0,\frac{1}{b}]^{2}.$ The IFS
$S_{\mathcal{A}}:=\left\{g_{(i,j)}\right\}_{(i,j)\in\mathcal{A}}$
is called a base-$b$ missing digit IFS. Its attractor $K$ is called a two-dimensional base $b$-missing digit set.
Note that it is direct to check that $S_{\mathcal{A}}$ satisfies the open set condition, with $O=(0,1)^{2}.$
Given $S_{\mathcal{A}}$ a base-$b$ missing digit IFS, we define for every $0\leq i\leq b-1,$
$p_{i,\mathcal{A}}=\frac{\#\left\{0\leq j\leq b-1:\ (j,i)\in\mathcal{A}\right\}}{\#\mathcal{A}}\in[0,1].$
Let $\mu_{0}$ be the Alfhors-regular self-similar measure on $K$, i.e. the self-similar measure solution to
$\mu_{0}(\cdot)=\sum_{(i,j)\in\mathcal{A}}\frac{1}{\#\mathcal{A}}g_{(i,j)}\mu_{0}(\cdot).$
It is easily seen that the orthogonal projection of $\mu_{0}$ on the $y$-axis, $\pi_{2}\mu_{0},$ is a self-similar measure associated with the IFS $\left\{g_{0,i}\right\}_{0\leq i\leq b-1}$ and the weights $(p_{0,i}=p_{i,\mathcal{A}})_{0\leq i\leq b-1}.$
Given $q\in\mathbb{R}$ and $0\leq i\leq b-1,$ we also set
$p_{i,\mathcal{A},q}=\frac{p_{i,\mathcal{A}}^{q}}{\sum_{0\leq j\leq b-1}p_{j,\mathcal{A}}^{q}}\text{ and }\kappa_{q,\mathcal{A}}=\frac{-\sum_{0\leq i\leq b-1}p_{i,\mathcal{A},q}\log p_{i,\mathcal{A}}}{\log b}.$
Regarding the random and dynamical covering by rectangles of $K$, our main result is the following.
Theorem 3.1.
Let $\frac{1}{s_{0}}\leq\tau_{1}\leq\tau_{2}$ be two real numbers and $S$ a base-$b$ two dimensional missing digit IFS. Let $(X_{n})_{n\in\mathbb{N}}$ be either an i.i.d. sequence of law $\mu_{0}$ or an orbit $(b^{n}x)_{n\in\mathbb{N}},$ where $x\in\mathbb{T}^{2},$ and
$W_{\tau_{1},\tau_{2}}=\limsup_{n\to+\infty}\Big(X_{n}+(-\frac{1}{n^{\tau_{1}}},\frac{1}{n^{\tau_{1}}})\times(-\frac{1}{n^{\tau_{2}}},\frac{1}{n^{\tau_{2}}})\Big).$
Write $\tau=\frac{\tau_{2}}{\tau_{1}}$ and define $v_{\tau}:\mbox{Spectr}(\pi_{2}(\mu_{0}))\to\mathbb{R}$ by
$v_{\tau}(\alpha)=s_{0}+(\tau-2)\alpha-(\tau-1)D_{\pi_{2}\mu_{0}}(\alpha).$
It is easily verified that $v_{\tau}$ is non increasing on $(-\infty,\kappa_{\frac{\tau-2}{\tau-1}}].$ Define, when possible (in particular when $\tau\neq 1$), $\beta_{\tau_{1},\tau_{2}}$ as the unique solution on $\infty,\kappa_{\frac{\tau-2}{\tau-1}}]$ to $v_{\tau}(\alpha)=\frac{1}{\tau_{1}}.$
Then, almost surely (or for $\mu_{0}$-almost every $x\in\mathbb{T}^{2}$):
(10) $\dim_{H}W_{\tau_{1},\tau_{2}}=\begin{cases}\frac{1}{\tau_{1}}\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \text{ if }\ \frac{1}{\tau_{1}}\leq\dim_{H}\mu_{0}-\dim_{H}\pi_{2}\mu_{0},\\ \frac{1}{\tau_{1}}-(\tau-1)(\beta_{\tau_{1},\tau_{2}}-D_{\pi_{2}\mu_{0}}(\beta_{\tau_{1},\tau_{2}}))\text{ if }\ \dim_{H}\mu_{0}-\dim_{H}\pi_{2}\mu_{0}\leq\frac{1}{\tau_{1}}\leq v_{\tau}(\kappa_{2}),\\ \frac{1+(\tau_{2}-\tau_{1})(s_{0}-2\kappa_{2}+D_{\pi_{2}\mu_{0}}(\kappa_{2})))}{\tau_{2}}\ \ \ \ \ \ \ \ \ \ \ \text{ if }\ \frac{1}{\tau_{1}}\geq v_{\tau}(\kappa_{2}),\end{cases}$
Remark 3.2.
  • (1)
    For simplicity, the results where formulate in the case $\tau_{1}\leq\tau_{2}$, but the results straightforwardly adapts to the case $\tau_{2}\leq\tau_{1}$ by switching the roles of $\tau_{1}$ and $\tau_{2}$ and considering $\pi_{1}\mu_{0}$ rather than $\pi_{2}\mu_{0}.$
  • (2)
    $\frac{1}{s_{0}}\leq\tau_{1}=\tau_{2},$ one recovers that $\dim_{H}W_{\tau_{1},\tau_{1}}=\frac{1}{\tau_{1}}.$
  • (3)
    When $S$ has uniform fibers, meaning that $p_{i,\mathcal{A}}=p_{j,\mathcal{A}}$ for every $0\leq i,j\leq b-1,$ then $\pi_{2}\mu_{0}$ is Alfhors-regular, which implies in particular that $\mbox{Spectr}(\pi_{2}\mu_{0})=\left\{\dim_{H}\pi_{2}\mu_{0}\right\}$ and $\dim_{H}\pi_{2}\mu_{0}=D_{\pi_{2}\mu_{0}}(\dim_{H}\pi_{2}\mu_{0}).$ Thus in this case, one obtains for every $\frac{1}{s_{0}}\leq\tau_{1}\leq\tau_{2},$
    $\dim_{H}W_{\tau_{1},\tau_{2}}=\min\left\{\frac{1}{\tau_{1}},\frac{1+(\tau_{2}-\tau_{1})(s_{0}-\dim_{H}\pi_{2}\mu_{0})}{\tau_{2}}\right\},$
    which is consistent with the sub-case where $\mathcal{A}=\mathcal{A}_{1}\times\mathcal{A}_{2},$ where $\mathcal{A}_{1},\mathcal{A}_{2}\subset\left\{0,...,b-1\right\}.$
The next section presents our result regarding the rectangular shrinking target problem.
3.2. Tree approximation
In this section, we study the `` tree approximation", which, as mentioned in introduction, can be seen as a reformulation of the classical shrinking target problem associated with the the mapping $T_{b}:\mathbb{T}^{2}\to\mathbb{T}^{2},$ defined by $T_{b}(x)=bx.$
Consider again $S=\left\{f_{1},..,f_{m}\right\}$, a two dimensional base-$b$ missing digit IFS of attractor $K$ and $\mu_{0}$ the Alfhors-regular self-similar measure on $K$. The following proposition is necessary to state our main result and will be established in the next section, as Proposition 4.1 applied with $\pi_{2}\nu$ (which can be identified with a $\times-b$ ergodic measure on $\mathbb{T}^{1}$).
Proposition 3.3.
Let $\nu$ be a $\times$-b (i.e. with respect to $T_{b}$) ergodic measure. Then, there exists $\alpha_{\nu}$ such that, for $\nu$-almost every $x$, one has
$\lim_{r\to 0^{+}}\frac{\log\pi_{2}\mu_{0}B\Big(\pi_{2}(x),r\Big)}{\log r}=\alpha_{\nu}.$
Theorem 3.4.
Let $\mu$ be a self-similar measure (with respect to $S$) and $\Lambda\subset\bigcup_{k\geq 1}\left\{1,...,m\right\}^{k}$ be a set of words such that
$\mu\Big(\limsup_{\underline{i}\in\Lambda}f_{\underline{i}}([0,1]^{2})\Big)=1.$
Then, for every $1\leq\tau_{1}\leq\tau_{2},$ for any $\times b$-ergodic measure $\nu$ with $\mathrm{supp}(\nu)\subset K$, writing
$V_{\tau_{1},\tau_{2}}(x)=\limsup_{\underline{i}\in\Lambda}\Big(f_{\underline{i}}(x)+(-|f_{\underline{i}}(K)|^{\tau_{1}},|f_{\underline{i}}(K)|^{\tau_{1}})\times(-|f_{\underline{i}}(K)|^{\tau_{2}},|f_{\underline{i}}(K)|^{\tau_{2}})\Big),$
for $\nu$-almost every $x$, one has
$\dim_{H}V_{\tau_{1},\tau_{2}}(x)\geq\min\left\{\frac{\dim_{H}\mu}{\tau_{1}},\frac{\dim_{H}\mu+(\tau_{2}-\tau_{1})(s_{0}-\alpha_{\nu})}{\tau_{2}}\right\}.$
Assume in addition that
$\lim_{|\underline{i}|\to+\infty}\frac{-\log\mu(f_{\underline{i}}([0,1]^{2}))}{|\underline{i}|\log b}=\dim_{H}\mu,$
then , for $\nu$-almost every $x$,
$\dim_{H}V_{\tau_{1},\tau_{2}}(x)=\min\left\{\frac{\dim_{H}\mu}{\tau_{1}},\frac{\dim_{H}\mu+(\tau_{2}-\tau_{1})(s_{0}-\alpha_{\nu})}{\tau_{2}}\right\}.$
Given $\nu$ a $\times b$-ergodic measure $\nu$ with $\mathrm{supp}(\nu)\subset K$, $\psi,\theta:\mathbb{N}\to\mathbb{R}_{+},$ define
$\begin{cases}\mathcal{N}_{1}=\left\{n:\ \psi(n)\geq\theta(n)\right\}\\ \mathcal{N}_{2}=\left\{n:\ \psi(n)<\theta(n)\right\}\end{cases}$
and $\beta_{\nu}$ the real number (which exists by Proposition 4.1) such that, for $\nu$-almost every $x$, one has
$\lim_{r\to 0^{+}}\frac{\log\pi_{1}\mu_{0}B\Big(\pi_{1}(x),r\Big)}{\log r}=\beta_{\nu}.$
Define also
$\limsup_{n\in\mathcal{N}_{1}}\min\left\{\frac{s_{0}}{\frac{\log\psi(n)}{-n\log b}},\frac{s_{0}+\Big(\frac{\log\theta(n)}{-n\log b}-\frac{\log\psi(n)}{-n\log b}\Big)(s_{0}-\alpha_{\nu})}{\frac{\log\theta(n)}{-n\log b}}\right\}$
$:=g_{1}(\psi,\theta,\nu)$
$\limsup_{n\in\mathcal{N}_{2}}\min\left\{\frac{s_{0}}{\frac{\log\theta(n)}{-n\log b}},\frac{s_{0}+\Big(\frac{\log\psi(n)}{-n\log b}-\frac{\log\theta(n)}{-n\log b}\Big)(s_{0}-\beta_{\nu})}{\frac{\log\psi(n)}{-n\log b}}\right\}$
$:=g_{2}(\psi,\theta,\nu)$
and, for $i=1,2,$ consider two non increasing sequences of integers $(n_{k,i})_{k\in\mathbb{N}}\subset\mathcal{N}_{i}^{\mathbb{N}}$ such that
$\lim_{k\to+\infty}\min\left\{\frac{s_{0}}{\frac{\log\psi(n_{k,1})}{-n_{k,1}\log b}},\frac{s_{0}+\Big(\frac{\log\theta(n_{k,1})}{-n_{k,1}\log b}-\frac{\log\psi(n_{k,1})}{-n_{k,1}\log b}\Big)(s_{0}-\alpha_{\nu})}{\frac{\log\theta(n_{k,1})}{-n_{k,1}\log b}}\right\}$
$:=g_{1}(\psi,\theta,\nu)$
and
$\lim_{k\to+\infty}\min\left\{\frac{s_{0}}{\frac{\log\theta(n_{k,2})}{-n_{k,2}\log b}},\frac{s_{0}+\Big(\frac{\log\psi(n_{k,2})}{-n_{k,2}\log b}-\frac{\log\theta(n_{k,2})}{-n_{k,2}\log b}\Big)(s_{0}-\beta_{\nu})}{\frac{\log\psi(n_{k,2})}{-n_{k,2}\log b}}\right\}$
$:=g_{2}(\psi,\theta,\nu).$
By applying 3.4 successively to $\Lambda_{1}=\bigcup_{k\geq 1}\left\{1,...,m\right\}^{n_{k,1}}$ and $\Lambda_{1}=\bigcup_{k\geq 1}\left\{1,...,m\right\}^{n_{k,2}}$ and $\mu_{0}$ , one obtains the following corollary.
Corollary 3.5.
Let $\psi,\theta:\mathbb{N}\to\mathbb{R}_{+}$ be two mappings such that
$\min\left\{\liminf_{n\to+\infty}\frac{\log\psi(n)}{-n\log b},\liminf_{n\to+\infty}\frac{\log\theta(n)}{-n\log b}\right\}\geq 1.$
Then
$\dim_{H}\limsup_{\underline{i}\in\bigcup_{k\geq 1}\left\{1,...,m\right\}^{k}}\Big(f_{\underline{i}}(x)+(-\psi(|\underline{i}|),\psi(|\underline{i}|)\times(-\theta(|\underline{i}|),\theta(|\underline{i}|))\Big)=\max\left\{g_{1}(\psi,\theta,\nu),g_{2}(\psi,\theta,\nu)\right\}.$
Remark 3.6.
Interestingly, unlike in the case of balls (i.e. for $\tau_{1}=\tau_{2}$), Theorem 3.4 shows that $\dim_{H}V_{\tau_{1},\tau_{2}}(x)$ depends in general on the choice of $x\in K.$ In particular, under these settings, the set of possible values (which are all attained) is
$\left\{\min\left\{\frac{\dim_{H}\mu}{\tau_{1}},\frac{\dim_{H}\mu+(\tau_{2}-\tau_{1})(s_{0}-\alpha)}{\tau_{2}}\right\}\ \alpha\in\mbox{Spectr}(\pi_{2}\mu_{0})\right\}.$
A careful reader will notice that the proof of Theorem 3.4 only requires $\pi_{2}\nu$ to be ergodic rather than $\nu.$ Thus the assumption of Theorem 3.4 can be weakened accordingly.
As mentioned in introduction, Corollary 3.5 was also obtained by Allen, Jordan and Ward in [2], using a different method, under the weaker assumption that the coding of $x=(x_{n}:=(x_{n}^{1},x_{n}^{2}))_{n\in\mathbb{N}}\in\mathcal{A}^{\mathbb{N}}$ satisfies that, writing $S=S_{\mathcal{A}},$ for each $0\leq j\leq b-1$ for which there exists $0\leq i\leq b-1$ so that $(i,j)\in\mathcal{A},$ one has
$\lim_{n\to+\infty}\frac{\#\left\{0\leq k\leq n:x_{n}^{2}=j\ \right\}}{n}=\kappa_{j},$
for some $0\leq\kappa_{j}<1.$
As a second application, we study the anisotropic approximation under digit frequency constraints. Such problems where for instance studied in [5, 8] To this end, we first recall a corollary of [5, Proposition 2.2]
Proposition 3.7 ([5]).
Let $(p_{1},...,p_{m})$ be a probability vector and
$\Lambda_{(p_{1},...,p_{m})}$
$=\left\{\underline{i}=(i_{1},...,i_{n}):\ \forall 1\leq k\leq m,\Big|\frac{1}{n}\#\left\{1\leq j\leq m:i_{j}=k\right\}-p_{k}\Big|\leq\sqrt{\frac{2\log\log n}{n}}\right\}.$
Then, if $\mu$ is the self-similar measure associated with $(p_{1},..,p_{m}),$ one has
$\mu\Big(\limsup_{\underline{i}\in\Lambda_{(p_{1},...,p_{m})}}f_{\underline{i}}\Big([0,1]^{2}\Big)\Big)=1.$
Our result related to approximation under digit frequencies is the following.
Corollary 3.8.
Let $(p_{1},...,p_{m})$ be a probability vector and
$\Lambda_{(p_{1},...,p_{m})}$
$=\left\{\underline{i}=(i_{1},...,i_{n}):\ \forall 1\leq k\leq m,\Big|\frac{1}{n}\#\left\{1\leq j\leq m:i_{j}=k\right\}-p_{k}\Big|\leq\sqrt{\frac{2\log\log n}{n}}\right\}.$
Then
$\dim_{H}\limsup_{\underline{i}\in\Lambda_{(p_{1},...,p_{m})}}\Big(f_{\underline{i}}(x)+(-|f_{\underline{i}}(K)|^{\tau_{1}},|f_{\underline{i}}(K)|^{\tau_{1}})\times(-|f_{\underline{i}}(K)|^{\tau_{2}},|f_{\underline{i}}(K)|^{\tau_{2}})\Big)$
$=\min\left\{\frac{\frac{-\sum_{1\leq i\leq m}p_{i}\log p_{i}}{\log b}}{\tau_{1}},\frac{\frac{-\sum_{1\leq i\leq m}p_{i}\log p_{i}}{\log b}+(\tau_{2}-\tau_{1})(s_{0}-\alpha_{\nu})}{\tau_{2}}\right\}.$
The next three sections are dedicated to the proofs of Theorem 3.1 and Theorem 3.4. The Section 4 establishes some useful general results regarding $\times-b$ ergodic measures, provides some recall on the mass transference principle for self-similar measures and establishes important content estimates, useful to the proof of both theorems. In the last section, we finally prove Theorem 3.1 and Theorem 3.4.
4. Some preliminaries to the proof of Theorems 3.1 and Theorem 3.4
4.1. Results regarding $\times-b$ ergodic measures
Let us fix $b\in\mathbb{N}$ and $\eta$ a $\times b$-ergodic measure on $\mathbb{T}$. For $0\leq i\leq b-1,$ let us denote $f_{i}$ the mapping defined for every $x\in\mathbb{T}$ by $f_{i}(x)=\frac{x+i}{b}$ and $S_{b}=\left\{f_{i}\right\}_{0\leq i\leq b-1}.$ The goal of this section is to prove the following result.
Proposition 4.1.
Let $\mu$ be a self-similar measure associated with the IFS $S_{b}$ and the probability vector $(p_{0},...,p_{b-1})$. Then for $\eta$-almost every $x$, one has
$\lim_{n\to+\infty}\frac{\log\mu\Big(D_{b,n}(x)\Big)}{-n\log b}=\lim_{r\to 0^{+}}\frac{\log\mu\Big(B(x,r)\Big)}{\log r}=\frac{-\sum_{0\leq i\leq b-1}\eta([\frac{i}{b},\frac{i+1}{b}])\log p_{i}}{\log b}.$
It is worth mentioning that equality between the left-handside and the right-handside can be obtained as a consequence of Birkhoff’s ergodic theorem. The difficulty here comes from the case where one considers a centered ball, as $\mu$ is not assume to be doubling. The proof of Proposition 4.1 relies on the following lemma.
Lemma 4.2.
Fix $\tau>1$ and write
$A_{b,\tau}=\left\{x:\ |x-\frac{k}{b^{n}}|\leq\frac{1}{b^{n\tau}}\text{ i.o. }\right\}.$
If $\eta$ has no atom then
$\eta\Big(A_{b,\tau}\Big)=0.$
Before proving Lemma 4.2, we show how it implies Proposition 4.1. Prior to that we start by a small classical lemma.
Lemma 4.3.
Let $T:\mathbb{R}^{d}\to\mathbb{R}^{d}$ be a measurable mapping and $\mu\in\mathcal{M}(\mathbb{R}^{d})$ a $T$-ergodic probability measure. Then $\mu$ is diffuse (i.e. has no atom) or carried by a periodic orbit (in which case it is purely atomic with finitely many atoms).
Proof.
Assume that there exists a measurable set $A$ with $\mu(A)>0$ and for any $x\in A,$ $\mu(\left\{x\right\})=0.$ Assume that there exists $y$ such that $\mu\Big(\left\{y\right\}\Big)>0.$ Then, for every $n\in\mathbb{N},$ as
$\left\{y\right\}\subset T^{-n}\Big(\left\{T^{n}(y)\right\}\Big),$
one has $\mu\Big(\left\{T^{n}(y)\right\}\Big)>0.$ In addition, by Poincarré’s recurrence Theorem, there exists $n$ such that $T^{n}(y)\in A,$ so that
$0=\mu\Big(\left\{T^{n}(y)\right\}\Big)>0.$
We conclude that there is no $y$ with $\mu\Big(\left\{y\right\}\Big)>0$ and $\mu$ is either purely atomic or diffuse. In the case where $\mu$ is purely atomic, from the fact that $\mu$ is an ergodic probability measure we deduce that $\mu$ must be carried by an eventually periodic orbit and from Poincarré’s recurrence Theorem that this eventually periodic orbit must be a periodic orbit. ∎
Proof.
First, recall that Since Proposition 4.1 is direct if $\eta$ is atomic, we assume now that $\eta$ has no atoms. For every $x\in\mathbb{T}$ and $0\leq i\leq b-1,$ write
$S_{N,i}(x)=\frac{1}{N}\sum_{n=0}^{N-1}\chi_{[\frac{i}{b},\frac{i+1}{b}[}(b^{n}x).$
By Birkhoff ergodic’s theorem, for $0\leq i\leq b-1,$ writing $\widetilde{p}_{i}=\eta([\frac{i}{b},\frac{i+1}{b}])$, for $\eta$-almost every $x$, for every $0\leq i\leq b-1,$
$S_{N,i}\to\widetilde{p}_{i}.$
Since $\mu$ is a Bernoulli measure, this implies that, for $\eta$-almost every $x$,
$\lim_{n\to+\infty}\frac{\log\mu\Big(D_{b,n}(x)\Big)}{-n\log b}=\frac{-\sum_{0\leq i\leq b-1}\widetilde{p_{i}}\log p_{i}}{\log b}:=\alpha.$
Let us fix $\tau>1.$ By Proposition 4.2, for $\eta$-almost every $x,$ there exists $N_{x}$ large enough so that for every $n\geq N_{x},$ for every $k\in\mathbb{N},$ one has
$|x-\frac{k}{b^{n}}|\geq\frac{1}{b^{n\tau}}.$
Thus we conclude that for every $n\geq N_{x},$ one has
$B(x,b^{-n\tau})\subset D_{b,n}(x)\subset B(x,b^{-n}).$
This yields
$\limsup_{r\to 0^{+}}\frac{\log\mu\Big(B(x,r)\Big)}{\log r}\leq\lim_{n\to+\infty}\frac{\log\mu\Big(D_{b,n}(x)\Big)}{-n\log b}\leq\tau\liminf_{r\to 0^{+}}\frac{\log\mu\Big(B(x,r)\Big)}{\log r}.$
Since $\tau>1$ was arbitrary, letting $\tau\to 1$ along a countable sequence proves the claim. ∎
We now prove Lemma 4.2.
Proof.
Fix $\tau>1$ and assume that
$\eta\Big(A_{b,\tau}\Big)>0.$
For every $0\leq i\leq b-1,$ for every $x\in\mathbb{T},$
$S_{N,i}(x)=\frac{1}{N}\sum_{n=0}^{N-1}\chi_{[\frac{i}{b},\frac{i+1}{b})}(b^{n}x)$
is an ergodic average. Thus writing $p_{i}=\eta\Big([\frac{i}{b},\frac{i+1}{b})\Big)$, for $\eta$-almost every $x\in A_{b,\tau},$ one has
$\lim_{N\to+\infty}S_{N,i}=p_{i}.$
On the other hand, for every $x\in A_{b,\tau}\setminus\left\{\frac{k}{b^{n}},n\in\mathbb{N},0\leq k\leq b^{n}\right\}$, there exists a sequence $(n_{k})_{m\in\mathbb{N}}$ of integers such that for every $k\in\mathbb{N},$
$|x-\frac{m_{k}}{b^{n_{k}}}|\leq\frac{1}{b^{\tau n_{k}}}$
for some $m_{k}\in\mathbb{N}.$ This, depending on the sign of $x-\frac{m_{k}}{b^{n_{k}}}$ implies that the coding on base $b$ of $x$ contains at least $(\tau-1)n_{k}$ 0 or b1b-1 in $n_{k}$-th position. Thus, for $j=0$ or b1,b-1, there exist a sub-sequence of integers $\widetilde{n}_{k}$ such that for every $k\in\mathbb{N},$ the expansion in base $b$ of $x$ contains $(\tau-1)\widetilde{n}_{k}$ $j$ in $\widetilde{n}_{k}$-th position. Assume now that there exists $1\leq i\neq j\leq b-1$ such that $p_{i}>0$ and fix $\varepsilon>0$ and $N$ large enough so that for every $n\geq N$,
$|S_{n,i}-p_{i}|\leq\varepsilon.$
Fix $\widetilde{n}_{k}>N.$ we have
$|S_{\widetilde{n}_{k},i}-p_{i}|\leq\varepsilon$
and the coding of $x$ contains $(\tau-1)\widetilde{n}_{k}$ $j\neq i$ in $\widetilde{n}_{k}$’th position. This yields that
$S_{\tau\widetilde{n}_{k}}\leq p_{i}+\varepsilon-\frac{\tau-1}{\tau}<p_{i}-\varepsilon$
provided that $\varepsilon$ was chosen small enough to begin with. We conclude that $p_{j}=1$ and $p_{i}=0$ for every $i\neq j.$
Recall that $x\in\mathbb{T}$ is called generic for $\eta$ if $\frac{1}{N}\sum_{k=0}^{N-1}\delta_{b^{k}x}$ converges weakly to $\eta$ and that $\eta$-almost every $x$ is generic. Thus for every interval $I\subset\mathbb{T}$, since $\eta(\partial I)=0,$ for any $\eta$-generic $x$, one has
(11) $\lim_{N\to+\infty}\frac{1}{N}\sum_{0\leq k\leq N-1}\chi_{I}(b^{n}x)=\eta(I).$
Now, fix $x$ $\eta$-generic and such that $S_{N,i}(x)\to p_{i}$ for every $0\leq i\leq b-1$ and $\underline{i}=(i_{1},...,i_{m})\in\left\{0,...,b-1\right\}^{m}$. Assume that there exists $1\leq k\leq m$ such that $i_{k}\neq j$, call $I_{\underline{i}}$ the projection in base $b$ of the cylinder $\underline{i}]$ on $\mathbb{T}.$ If $\eta(I_{\underline{i}})>0,$ by (11), writing $(x_{n})_{n\in\mathbb{N}}$ the coding of $x$ in base $b$, one gets
$\lim_{N\to+\infty}\frac{1}{N}\sum_{0\leq n\leq N-1}\chi_{I_{\underline{i}}}(b^{n}x)=\lim_{N\to+\infty}\frac{1}{N}\sum_{0\leq n\leq N-1}\chi_{\underline{i}}((x_{n},...,x_{n+m-1}))=\eta(I_{\underline{i}}).$
This implies in particular that
$\liminf_{N\to+\infty}\frac{1}{N}\sum_{0\leq n\leq N-1}\chi_{[\frac{i_{k}}{b},\frac{i_{k}+1}{b}[}(b^{n}x)=\liminf_{N\to+\infty}\frac{1}{N}\sum_{0\leq n\leq N-1}\chi_{i_{k}}(x_{n})\geq\frac{\eta(I_{\underline{i}})}{2m}>0,$
which is a contradiction. Thus $\eta([\underline{i}])=0$ whenever there exists $i_{k}$ such that $i_{k}\neq j$ but this implies $\eta=\delta_{0}$. The assumption $\eta\Big(A_{b,\tau}\Big)>0$ was therefore false. ∎
We also isolate the following corollary, which will be useful later on.
Corollary 4.4.
Let $\nu\in\mathcal{M}(\mathbb{T}^{1})$ be a non atomic $\times$ b ergodic measure. Then for every $0<\varepsilon<1,$ for $\eta$-almost every $x\in\mathbb{T}^{1},$ there exists $n_{x,\varepsilon}>0$ such that for every $n\geq n_{x,\varepsilon},$ one has
$D_{b,n}(x)\subset B(x,b^{-n})\subset D_{b,\lfloor(1-\varepsilon)n\rfloor}(x)$
4.2. Recall on mass transference principle for self-similar measures
This section is dedicated to some recall on a key result one will use in order to prove Theorem 3.1 and Theorem 3.4.
Theorem 4.5.
Let $\mu\in\mathcal{M}(\mathbb{R}^{d})$ be a self-similar measure. Let $(B_{n})_{n\in\mathbb{N}}$ be a sequence of balls such that $|B_{n}|\to 0$ and
$\mu\Big(\limsup_{n\to+\infty}B_{n}\Big)=1.$
Let $0\leq s<\dim_{H}\mu$ and $(U_{n})_{n\in\mathbb{N}}$ be a sequence of open sets such that for every $n\in\mathbb{N}$,
  • $U_{n}\subset B_{n},$
  • $\mathcal{H}^{s}_{\infty}\Big(U_{n}\cap\mathrm{supp}(\mu)\Big)\geq\mu\Big(B_{n}\Big).$
Then $\dim_{H}\limsup_{n\to+\infty}U_{n}\geq s.$
4.3. Content estimates
Let us fix $S$ a $b$-adic IFS, call $\mu_{0}$ the corresponding Alfhors-regular self-similar measure and $\widetilde{K}$ its attractor. In this section, in order to avoid possible confusion between $b$-adic intervals and $b$-adic cubes of $0,1]^{2},$ we will denote II $b$-adic intervals and $D$ $b$-adic cubes. In addition, we will identify when necessary the measure $\pi_{2}\mu_{0}$ with its one dimensional natural counterpart.
Let us start by remarking that, if there exists a $b$-adic cube $D$ such that $\mu_{0}(\partial D)>0,$ then either $\pi_{1}\mu_{0}$ or $\pi_{2}\mu_{0}$ has an atom. Since these measures are self-similar, one obtains that this measure is a Dirac mass so that $\mu_{0}$ is supported on an horizontal or vertical line. In the rest of the section, we assume that $\mu_{0}$ is not supported on an horizontal or vertical line, hence satisfies that $\mu_{0}(\partial D)=0$ for any $b$-adic cube $D$.
Given $\eta\in\mathcal{M}(\mathbb{R}^{d})$ a Borel set $A\in\mathcal{M}(\mathbb{R}^{d})$ and $s\geq 0,$ define
$\mathcal{H}^{\eta,s}_{\infty}(A)=\inf\left\{\mathcal{H}^{s}_{\infty}(E),\ E\subset A:\ \mu(E)=\mu(A)\right\}.$
The following result, established as [9, Theorem 2.6].
Theorem 4.6.
Let $\Omega$ be an open set and $0\leq s\leq\dim_{H}\mu_{0}:=s_{0}$. Then there exists a constant $C$ (depending only on $\mu_{0}$ and $s$) such that
[…]
Write
$F=\bigcup_{n\geq 0,D\in\mathcal{D}_{n}}\partial D.$
As a consequence of Theorem 4.6 and the fact that $\mu_{0}(F)=0,$ there exists a constant $C>0$ such that for any open sets $\Omega$
$\mathcal{H}^{s}_{\infty}(\Omega\cap K)\leq\mathcal{H}^{s}_{\infty}(\Omega\cap\widetilde{K})\leq C\mathcal{H}^{s}_{\infty}(\Omega\cap K),$
where
$K=\widetilde{K}\setminus F.$
Moreover, since $S$ satisfies the open set condition with open set $(0,1)^{2}$, it is direct to verify that for any open set $\Omega\subset[0,1]^{2}$ and any $b$-adic cube $D$, one has
$f_{D}(\Omega\cap K)=f_{D}(\Omega)\cap K,$
where $f_{D}$ denotes the canonical contraction from $0,1)^{2}$ to $D$.
Proposition 4.7.
Let $I$ be a $b$-adic interval, $x\in[0,1]\times I\cap K$ and let us write
$\mbox{Str}(I)=[0,1]\times I.$
Let $K$ be a base $b$ missing digit set. Then for every $0<s\leq\dim_{H}K$, one has
$\frac{C_{s}}{-\log|I|}\leq\frac{\mathcal{H}^{s}_{\infty}\Big(\mbox{Str}(I)\cap K\Big)}{\min_{0\leq\gamma\leq 1}b^{-(s-s_{0})\lfloor\gamma n\rfloor}\times\pi_{2}\mu_{0}(D_{\lfloor\gamma n\rfloor}(\pi_{2}(x)))}\leq 1$
Proof.
Notice that, because $\pi_{2}$ induces a semi-conjugacy between $S$ and the projected IFS on $\left\{0\right\}\times[0,1]$ (with the base $b$ shifts), denoting $\sigma$ both of these shifts, one has
$\begin{cases}\mu_{0}(\mbox{Str}(I))=\pi_{2}\mu_{0}(D_{n}(\pi_{2}(x)))\\ \mu_{0}\Big(\mbox{Str}\Big(I_{\lfloor(1-\gamma)n\rfloor}(\pi_{2}(\sigma^{\lfloor\gamma n\rfloor}(x)))\Big)\Big)=\pi_{2}\mu_{0}\Big(D_{\lfloor(1-\gamma)n\rfloor}(\sigma^{\lfloor\gamma n\rfloor}(\pi_{2}(x)))\Big)\end{cases}.$
In addition, denoting $(p_{i})_{1\leq i\leq b}$ the probability vector associated with the self-similar measure $\pi_{2}\mu_{0}$ and $(y_{n})_{n\in\mathbb{N}}$ the coding in base $b$ of $\pi_{2}(x),$ one has
$\begin{cases}\pi_{2}\mu_{0}(D_{n}(\pi_{2}(x)))=\prod_{k=1}^{n}p_{y_{i}}\\ \pi_{2}\mu_{0}\Big(D_{\lfloor(1-\gamma)n\rfloor}(\sigma^{\lfloor\gamma n\rfloor}(\pi_{2}(x)))\Big)=\prod_{k=\lfloor\gamma n\rfloor+1}^{\lfloor\gamma n\rfloor+\lfloor(1-\gamma)n\rfloor}p_{y_{i}}.\end{cases}$
Thus there exists a constant $C>0$ such that for any $0\leq\gamma\leq 1,$
$C^{-1}\prod_{i=1}^{\lfloor\gamma n\rfloor}p_{y_{i}}\leq\frac{\mu_{0}\Big(\mbox{Str}(I)\Big)}{\mu_{0}\Big(\mbox{Str}\Big(I_{\lfloor(1-\gamma)n\rfloor}(\pi_{2}(\sigma^{\lfloor\gamma n\rfloor}(x)))\Big)\Big)}\leq C\prod_{i=1}^{\lfloor\gamma n\rfloor}p_{y_{i}}=C\pi_{2}\mu_{0}(D_{\lfloor\gamma n\rfloor}(\pi_{2}(x))).$
Thus it is sufficient to show that
$\frac{C_{s}}{-\log|I|}\leq\frac{\mathcal{H}^{s}_{\infty}\Big(\mbox{Str}(I)\cap K\Big)}{\min_{0\leq\gamma\leq 1}b^{-(s-s_{0})\lfloor\gamma n\rfloor}\times\frac{\mu_{0}\Big(\mbox{Str}(I)\Big)}{\mu_{0}\Big(\mbox{Str}\Big(I_{\lfloor(1-\gamma)n\rfloor}(\pi_{2}(\sigma^{\lfloor\gamma n\rfloor}(x)))\Big)\Big)}}\leq 1.$
We first establish the lower-bound. Let us first notice that
$\mbox{Str}(I)=\bigcup_{D\in\mathcal{D}_{b,n}:D\cap\mbox{Str}(I)\neq\emptyset}D\cap K.$
In addition, for any $D\in\mathcal{D}_{b},$ since $f_{D}$ is homothetic of Lipshitz constant $|D|,$ one has
$\mathcal{H}^{s}_{\infty}(D\cap K)=\mathcal{H}^{s}_{\infty}(f_{D}(K))=|D|^{s}\mathcal{H}^{s}_{\infty}(K)=\kappa_{s}|D|^{s}.$
Thus, up to multiplying by some constant depending on $s$ if one must, we may consider only coverings by $b$-adic cubes of generation smaller than $n$.
Fix a covering $\mathcal{C}$ of $\mbox{Str}(I)$ by $b$-adic cubes of generation smaller than $n.$ There must exist $0\leq k\leq n$ such that
$\mu_{0}\Big(\bigcup_{D\in\mathcal{C}:\ \mbox{gen}(D)=k}D\cap\mbox{Str(I)}\Big)\geq\frac{1}{n+1}\mu_{0}(R).$
In addition, notice that, for every $D_{1},D_{2}\in\left\{D\in\mathcal{C}:\ \mbox{gen}(D)=k\right\}$, one has
$f^{-1}_{D_{1}}(\mbox{Str}(I))=f^{-1}_{D_{2}}(\mbox{Str}(I)),$
so that, by self-similarity,
$\mu_{0}(D_{1}\cap\mbox{Str}(I))$ $=\mu_{0}(D_{1})\times\mu_{0}\Big(f^{-1}_{D_{1}}(\mbox{Str}(I)\Big)$
$=b^{-ks_{0}}\mu_{0}\Big(f^{-1}_{D_{1}}(\mbox{Str}(I)\Big)=\mu_{0}(D_{2})\times\mu_{0}\Big(f^{-1}_{D_{2}}(\mbox{Str}(I)\Big)=\mu_{0}(D_{2}\cap\mbox{Str}(I)).$
Moreover, fixing $x\in D_{1}\cap K,$ one has
$\mu_{0}(D_{1}\cap\mbox{Str}(I))=b^{-ks_{0}}\mu_{0}\Big(\mbox{Str}\Big(I_{n-k}(\pi_{2}(\sigma^{k}(x)))\Big)\Big),$
which implies that
$\left\{D\in\mathcal{C}:\ \mbox{gen}(D)=k\right\}\geq\frac{\mu_{0}(R)}{(n+1)b^{-ks_{0}}\mu_{0}\Big(\mbox{Str}\Big(I_{n-k}(\pi_{2}(\sigma^{k}(x)))\Big)\Big)}.$
Thus
$\sum_{D\in\mathcal{C}}|D|^{s}$ $\geq\sum_{D\in\mathcal{C}:\ \mbox{gen}(D)=k}|D|^{s}=b^{-ks}\#\left\{D\in\mathcal{C}:\ \mbox{gen}(D)=k\right\}$
$\geq b^{-k(s-s_{0})}\times\frac{1}{n+1}\times\frac{\mu_{0}(R)}{\mu_{0}\Big(\mbox{Str}\Big(I_{n-k}(\pi_{2}(\sigma^{k}(x)))\Big)\Big)}.$
Noticing that $n+1\leq 2n=-2\log_{b}|I|$ and that the above inequality holds up to some multiplicative constants depending on $s$, the lower-bound is obtained by setting $k=\lfloor n\gamma\rfloor$, with $\frac{k}{n}\leq\gamma\leq\frac{k+1}{n}.$
The upper-bound is simply obtained by covering at the scale $k=\lfloor\gamma n\rfloor$ corresponding to the minimum and applying the same estimates. ∎
Given $x\in K,$ Proposition 4.7, actually allows one to estimate $\mathcal{H}^{s}_{\infty}(R\cap K)$ for any rectangle of the form $J\times I$, where $I$ is $b$-adic interval and $J$ is any interval centered on $x$.
Corollary 4.8.
Let $I$ be a $b$-adic interval and $J$ an interval such that $\frac{1}{b}J\cap K\neq\emptyset$ and $\frac{|J|}{b}\geq|I|.$ Let $D$ be the largest b-adic interval contained in $J$. Notice that $f_{D}^{-1}(R)$ is a stripe of the form $0,1]\times\widetilde{I}$, where $\widetilde{I}$ is a $b$-adic interval. Moreover one has
$b^{-s\mbox{gen}(D)}\mathcal{H}^{s}_{\infty}(f_{D}^{-1}(R)\cap K)\leq\mathcal{H}^{s}_{\infty}(K\cap R)\leq 3\times b^{-s\mbox{gen}(D)}\mathcal{H}^{s}_{\infty}(f_{D}^{-1}(R)\cap K)$
Proof.
Simply notice that $D\cap R$ intersects $K$ and contains a full stripe. In addition, $R$ intersects at most two more dyadic cubes of the same generation as $D$, say $D_{1}$ and $D_{2}$, but by self-similarity
$\max_{i=1,2}\mathcal{H}^{s}_{\infty}(D_{i}\cap R\cap K)\leq\mathcal{H}^{s}_{\infty}(R\cap D\cap K)$
so that
$\mathcal{H}^{s}_{\infty}(R\cap D\cap K)\leq\mathcal{H}^{s}_{\infty}(K\cap R)\leq 3\mathcal{H}^{s}_{\infty}(R\cap D\cap K).$
In addition, since $f_{D}$ is homotethic,
$\mathcal{H}^{s}_{\infty}(f_{D}^{-1}(R)\cap K)=|D|^{-s}\mathcal{H}^{s}_{\infty}(D\times I\cap K).$
Proposition 4.9.
Let $\nu\in\mathcal{M}([0,1]^{2})$ be a $\times$-b ergodic measure. Then, for $\nu$-almost every $x$ for every $\varepsilon>0$, there exists $r_{x}>0$ such that for every $r\leq r_{x}$, for every $0\leq s\leq s_{0},$ one has
$r^{\max\left\{s\tau_{1},s\tau_{2}-(\tau_{2}-\tau_{1})(s_{0}-\alpha_{\nu})\right\}+\varepsilon}\leq\mathcal{H}^{s}_{\infty}\Big((x+(-r^{\tau_{1}},r^{\tau_{1}})\times(-r^{\tau_{2}},r^{\tau_{2}}))\cap K\Big)\leq r^{\max\left\{s\tau_{1},s\tau_{2}-(\tau_{2}-\tau_{1})(s_{0}-\alpha_{\nu})\right\}-\varepsilon}$
Proof.
Fix $\varepsilon_{0}>0.$ Recall that, since $\pi_{2}\nu$ is $\times-b$ ergodic, either $\nu$ is a periodic measure (hence is carried by finitely many atoms) or $\pi_{2}\nu$ is diffuse, which case, by Corollary 4.4, there exists $r_{1}>0$ small enough so that for every $r\leq r_{1},$ one has
$D_{\lfloor\frac{-\tau_{1}\log r}{\log b}\rfloor}(\pi_{2}(x))\subset B\Big(\pi_{2}(x),r^{\tau_{1}}\Big)\subset D_{\lfloor\frac{-\tau_{1}(1-\varepsilon_{0})\log r}{\log b}\rfloor}(\pi_{2}(x)).$
Thus, writing $\tau=\frac{\tau_{2}}{\tau_{1}},$ and
$R_{n,\tau}(x)=(\pi_{1}(x)-b^{n},\pi_{1}(x)+b^{n})\times I_{\lfloor n\tau\rfloor}(\pi_{2}(x)),$
it is enough to show that for every large enough $n$, one has
$b^{-n\max\left\{s,s\tau-(\tau-1)(s_{0}-\alpha_{\nu})\right\}+\varepsilon}\leq\mathcal{H}^{s}_{\infty}\Big(R_{n,\tau}(x)\cap K\Big)\leq b^{-n\max\left\{s,s\tau-(\tau-1)(s_{0}-\alpha_{\nu})\right\}-\varepsilon}.$
Notice that $R_{n,\tau}(x)$ can be covered by at most $3$, from left to right, say $D_{1},D_{2}:=D_{n}(x),D_{3}$ $b$-adic cubes of generation $n$. Writing $f_{D_{1}},f_{D_{2}},f_{D_{3}}$ the corresponding contractions, one has
$f_{D_{1}}^{-1}(R_{n,\tau}(x)),f_{D_{3}}^{-1}(R_{n,\tau}(x))\subset f_{D_{2}}^{-1}(R_{n,\tau}(x)).$
This yields that
$\mathcal{H}^{s}_{\infty}\Big(R_{n,\tau}(x)\cap K\cap D_{n}(x)\Big)\leq\mathcal{H}^{s}_{\infty}\Big(R_{n,\tau}(x)\Big)\leq 3\mathcal{H}^{s}_{\infty}\Big(R_{n,\tau}(x)\cap K\cap D_{n}(x)\Big).$
Also,
$R_{n,\tau}(x)\cap D_{n}(x)=I_{n}(\pi_{1}(x))\times I_{\lfloor n\tau\rfloor}(\pi_{2}(x)).$
By Proposition 4.1, there exists $n_{1}\in\mathbb{N}$ such that, setting
$A_{n_{1}}=\left\{y\in[0,1]^{2}:\ \forall k\geq n_{1}\ b^{-k(\alpha_{\nu}+\varepsilon)}\leq\pi_{2}\mu_{0}(I_{k}(\pi_{2}(y)))\leq b^{-k(\alpha_{\nu}-\varepsilon)}\right\}$
satisfies that
$\nu(A_{n_{1}})\geq 1-\varepsilon_{0}.$
By ergodicity, for $\nu$-almost every $x$, there exists $n_{3}$ large enough so that for every $k\geq n_{3},$ there exists $(1-2\varepsilon_{0})k\leq k_{1}\leq k\leq k_{2}\leq(1+2\varepsilon_{0})k$ for which $b^{k_{i}}x\in A_{n_{1}},$ for $i=1,2.$ Thus provided that $n$ is large enough, fix $(1-2\varepsilon_{0})n\leq n_{1}\leq n\leq n_{2}\leq(1+2\varepsilon_{0})n$ and remark that
$I_{n_{2}}(\pi_{1}(x))\times I_{\lfloor n\tau\rfloor}(\pi_{2}(x))\subset I_{n}(\pi_{1}(x))\times I_{\lfloor n\tau\rfloor}(\pi_{2}(x))\subset I_{n_{1}}(\pi_{1}(x))\times I_{\lfloor n\tau\rfloor}(\pi_{2}(x)).$
Since $f_{D_{n_{1}}(x)},f_{D_{n_{2}}(x)}$ are homothetic, we have
$\mathcal{H}^{s}_{\infty}\Big(f_{D_{n_{i}}(x)}^{-1}\Big(I_{n_{i}}(\pi_{1}(x))\times I_{\lfloor n\tau\rfloor}(\pi_{2}(x))\cap K\Big)\Big)=|D_{i}|^{-s}\mathcal{H}^{s}_{\infty}\Big(I_{n_{i}}(\pi_{1}(x))\times I_{\lfloor n\tau\rfloor}(\pi_{2}(x))\cap K\Big).$
Moreover, remark that
$f_{D_{n_{i}}(x)}^{-1}\Big(I_{n_{i}}(\pi_{1}(x))\times I_{\lfloor n\tau\rfloor}(\pi_{2}(x))\Big)=[0,1]\times I_{\lfloor n\tau\rfloor-n_{i}}(\pi_{2}(b^{n_{i}}(x))).$
Applying Proposition 4.7 yields
$\frac{C}{\log(n)}\leq\frac{\mathcal{H}^{s}_{\infty}\Big(f_{D_{n_{i}}(x)}^{-1}\Big(I_{n_{i}}(\pi_{1}(x))\times I_{\lfloor n\tau\rfloor}(\pi_{2}(x))\Big)\cap K\Big)}{\min_{0\leq\gamma\leq 1}b^{-(s-s_{0})\lfloor\gamma(\lfloor n\tau\rfloor-n_{i})\rfloor}\times\pi_{2}\mu_{0}(I_{\lfloor\gamma(\lfloor n\tau\rfloor-n_{i})\rfloor}(\pi_{2}(b^{n_{i}}(x))))}\leq 1.$
Noticing that $(\tau-1-2\varepsilon_{0})n\leq n\tau-n_{i}\leq(\tau-1+2\varepsilon_{0})n$, one has
$\begin{cases}\frac{\min_{0\leq\gamma\leq 1}b^{-(s-s_{0})\lfloor\gamma(\lfloor n\tau\rfloor-n_{i})\rfloor}\times\pi_{2}\mu_{0}(I_{\lfloor\gamma(\lfloor n\tau\rfloor-n_{i})\rfloor}(\pi_{2}(b^{n_{i}}(x))))}{\min_{k\geq(\tau-1-2\varepsilon_{0})n}b^{-(s-s_{0})k}\times\pi_{2}\mu_{0}(D_{k}(\pi_{2}(b^{n_{i}}(x))))}\leq 1\\ \frac{\min_{0\leq\gamma\leq 1}b^{-(s-s_{0})\lfloor\gamma(\lfloor n\tau\rfloor-n_{i})\rfloor}\times\pi_{2}\mu_{0}(I_{\lfloor\gamma(\lfloor n\tau\rfloor-n_{i})\rfloor}(\pi_{2}(b^{n_{i}}(x))))}{\min_{k\geq(\tau-1+2\varepsilon_{0})n}b^{-(s-s_{0})k}\times\pi_{2}\mu_{0}(I_{k}(\pi_{2}(b^{n_{i}}(x))))}\geq 1.\end{cases}$
This yields that, for some $\kappa>0,$
$b^{-\kappa\varepsilon_{0}n}\leq\frac{\min_{0\leq\gamma\leq 1}b^{-(s-s_{0})\lfloor\gamma(\lfloor n\tau\rfloor-n_{i})\rfloor}\times\pi_{2}\mu_{0}(D_{\lfloor\gamma(\lfloor n\tau\rfloor-n_{i})\rfloor}(\pi_{2}(b^{n_{i}}(x))))}{\min_{k\geq(\tau-1)n}b^{-(s-s_{0})k}\times\pi_{2}\mu_{0}(I_{k}(\pi_{2}(b^{n_{i}}(x))))}\leq b^{\kappa\varepsilon_{0}n}.$
There exists a constant $\beta_{n_{1}}$, depending only on $n_{1}$, such that such that
$\beta_{n_{1}}\leq\min_{k\geq n_{1}}b^{-(s-s_{0})k}\times\pi_{2}\mu_{0}(I_{k}(\pi_{2}(b^{n_{i}}(x))))\leq 1$
and, since $b^{n_{i}}(x)\in A_{n_{1}},$ for every $(\tau-1)n\leq k\leq n_{1},$ one has
$b^{-k(\alpha_{\nu}+\varepsilon_{0})}\leq\pi_{2}\mu_{0}(I_{k}(\pi_{2}(b^{n_{i}}(x))))\leq b^{-k(\alpha_{\nu}-\varepsilon_{0})}.$
This yields
$\min_{(\tau-1)n\leq k\leq n_{1}}b^{-k(s-s_{0}+\alpha_{\nu}+\varepsilon_{0})}$ $\leq\min_{(\tau-1)n\leq k\leq n_{1}}b^{-(s-s_{0})k}\times\pi_{2}\mu_{0}(I_{k}(\pi_{2}(b^{n_{i}}(x))))$
$\min_{(\tau-1)n\leq k\leq n_{1}}b^{-k(s-s_{0}+\alpha_{\nu}-\varepsilon_{0})}$ $\geq\min_{(\tau-1)n\leq k\leq n_{1}}b^{-(s-s_{0})k}\times\pi_{2}\mu_{0}(I_{k}(\pi_{2}(b^{n_{i}}(x)))).$
Hence, for every $s\leq s_{0}-\alpha_{\nu}-\varepsilon_{0}$ or $s\geq s_{0}-\alpha_{\nu}-\varepsilon_{0}$ since, $k\mapsto k(s-s_{0}+\alpha_{\nu}-\varepsilon_{0})$ and $k\mapsto k(s-s_{0}+\alpha_{\nu}+\varepsilon_{0})$ are monotonic the infimimum is obtained for $k=\lfloor(\tau-1)n\rfloor+1$ or $k=n_{1}$ and there exists a constant $\theta_{n_{1}}$ such that
$\min\left\{\beta_{n_{1}},b^{-(\tau-1)n(s-s_{0}+\alpha_{\nu}+\varepsilon_{0})}\right\}$ $\leq\min_{(\tau-1)n\leq k\leq n_{1}}b^{-(s-s_{0})k}\times\pi_{2}\mu_{0}(I_{k}(\pi_{2}(b^{n_{i}}(x))))$
$\min_{(\tau-1)n\leq k\leq n_{1}}b^{-(s-s_{0})k}\times\pi_{2}\mu_{0}(I_{k}(\pi_{2}(b^{n_{i}}(x))))$ $\leq\min\left\{\beta_{n_{1}},b^{-(\tau-1)n(s-s_{0}+\alpha_{\nu}-\varepsilon_{0})}\right\}.$
Finally, since $s\mapsto\mathcal{H}^{s}_{\infty}(\cdot)$ is non increasing, for every $s\leq s_{0},$ (in particular for $s_{0}-\alpha_{\nu}-\varepsilon_{0}\leq s\leq s_{0}-\alpha_{\nu}+\varepsilon_{0}$), there exists $\omega_{n_{1}}>0$ such that one has
$\frac{C}{\log n}\min\left\{\omega_{n_{1}},b^{-(\tau-1)n(s-s_{0}+\alpha_{\nu}+3\varepsilon_{0})}\right\}$ $\leq\mathcal{H}^{s}_{\infty}\Big(f_{D_{n_{i}}(x)}^{-1}\Big(I_{n_{i}}(\pi_{1}(x))\times I_{\lfloor n\tau\rfloor}(\pi_{2}(x))\Big)\cap K\Big)$
$\min\left\{\omega_{n_{1}},b^{-(\tau-1)n(s-s_{0}+\alpha_{\nu}-3\varepsilon_{0})}\right\}$ $\geq\mathcal{H}^{s}_{\infty}\Big(f_{D_{n_{i}}(x)}^{-1}\Big(I_{n_{i}}(\pi_{1}(x))\times I_{\lfloor n\tau\rfloor}(\pi_{2}(x))\Big)\cap K\Big),$
which implies that
$\frac{C}{\log n}\min\left\{b^{-(1+\varepsilon_{0})ns}\omega_{n_{1}},b^{-(\tau-1)n(s-s_{0}+\alpha_{\nu}+3\varepsilon_{0})-(1+\varepsilon_{0})ns}\right\}$ $\leq\mathcal{H}^{s}_{\infty}\Big(I_{n_{i}}(\pi_{1}(x))\times I_{\lfloor n\tau\rfloor}(\pi_{2}(x))\cap K\Big)$
$\min\left\{b^{-(1-\varepsilon_{0})ns}\omega_{n_{1}},b^{-(\tau-1)n(s-s_{0}+\alpha_{\nu}+3\varepsilon_{0})-(1-\varepsilon_{0})ns}\right\}$ $\geq\mathcal{H}^{s}_{\infty}\Big(I_{n_{i}}(\pi_{1}(x))\times I_{\lfloor n\tau\rfloor}(\pi_{2}(x))\cap K\Big).$
Recalling the inclusion and since $\log(n)=b^{o(n)}$, provided that $n$ is large enough, one gets
$\min\left\{b^{-(1+4\varepsilon_{0})ns},b^{-(\tau-1)n(s-s_{0}+\alpha_{\nu}+4\varepsilon_{0})-(1+4\varepsilon_{0})ns}\right\}$ $\leq\mathcal{H}^{s}_{\infty}(R_{n,\tau}(x)\cap K)$
$\min\left\{b^{-(1-4\varepsilon_{0})ns},b^{-(\tau-1)n(s-s_{0}+\alpha_{\nu}-4\varepsilon_{0})-(1-4\varepsilon_{0})ns}\right\}$ $\geq\mathcal{H}^{s}_{\infty}(R_{n,\tau}(x)\cap K).$
Moreover
$\min\left\{b^{-(1+4\varepsilon_{0})ns},b^{-(\tau-1)n(s-s_{0}+\alpha_{\nu}+4\varepsilon_{0})-(1+4\varepsilon_{0})ns}\right\}$ $\geq b^{-n\max\left\{s,s\tau-(\tau-1)(s_{0}-\alpha)\right\}+(4s_{0}+4\tau)\varepsilon_{0}}$
$\min\left\{b^{-(1-4\varepsilon_{0})ns},b^{-(\tau-1)n(s-s_{0}+\alpha_{\nu}-4\varepsilon_{0})-(1-4\varepsilon_{0})ns}\right\}$ $\leq b^{-n\max\left\{s,s\tau-(\tau-1)(s_{0}-\alpha)\right\}-(4s_{0}+4\tau)\varepsilon_{0}}.$
One concludes by taking $\varepsilon_{0}$ so small that $(4s_{0}+4\tau)\varepsilon_{0}\leq\varepsilon.$
5. Proof of Theorem 3.4 and Theorem 3.1
5.1. Proof of Theorem 3.4
In this section again, given $x\in[0,1]$ and $n\in\mathbb{N}$, we will write $I_{n}(x)$ the $b$-adic interval of generation $n$ containing $x$ and given $y\in[0,1]^{2},$ $D_{n}(y)$ still denotes the $b$-adic cube of generation $n$ containing $y$. We fix $S$ , $\mu$, $1\leq\tau_{1}<\tau_{2}$ as in Theorem 3.4 and write $\tau=\frac{\tau_{2}}{\tau_{1}}$.
We start by dealing with the case where $\mu_{0}$ is supported on an horizontal or vertical line.
Assume that $\mu_{0}$ is supported on an horizontal line. To prove that Theorem 3.4 provides the correct estimate, one simply needs to prove that one recovers the same results as the case of of balls i.e. (see [7]) that
$\dim_{H}V_{\tau_{1},\tau_{2}}\geq\frac{\dim_{H}\mu}{\tau_{1}}.$
In this case $\pi_{2}\mu_{0}$ is an atom so that for any $\nu$ ergodic, $\alpha_{\nu}=0.$ Moreover
$\frac{\dim_{H}\mu}{\tau_{1}}$ $\leq\frac{\dim_{H}\mu+(\tau_{2}-\tau_{1})s_{0}}{\tau_{2}}$
$\Leftrightarrow\frac{\dim_{H}\mu}{\tau_{1}}$ $\leq\frac{\dim_{H}\mu}{\tau\tau_{1}}+(1-\frac{1}{\tau})s_{0}$
$\frac{\dim_{H}\mu}{\tau_{1}}$ $\leq s_{0}$
which is always satisfied as $\tau_{1}\geq\frac{1}{s_{0}}$ and $\dim_{H}\mu\leq 1.$ Thus one recovers the correct estimate in that case.
The case where $\mu_{0}$ is carried by a vertical line is straightforward as $\dim_{H}\pi\mu_{0}=s_{0}$ and $\frac{1}{\tau_{2}}\leq\frac{1}{\tau_{1}}.$
We now assume that $\mu_{0}$ is not carried by an horizontal or vertical line and we use the notations of Subsection 4.3. Before proving Theorem 3.4, we establish some estimates. Let us fix $\nu$ an ergodic measure with respect to $S$ and $\alpha_{\nu}$ such that for $\nu$-almost every $x$,
$\dim(\pi_{2}(x),\pi_{2}\mu_{0})=\alpha_{\nu}.$
In what follows, given a word $\underline{i}=(i_{1},...,i_{n})\in\left\{1,...,m\right\}^{n},$ we will write
$\begin{cases}R_{\underline{i}}(x)=f_{\underline{i}}(x)+(-|f_{\underline{i}}(K)|^{\tau_{1}},|f_{\underline{i}}(K)|^{\tau_{1}})\times(-|f_{\underline{i}}(K)|^{\tau_{2}},|f_{\underline{i}}(K)|^{\tau_{2}})\\ D_{\underline{i}}=f_{\underline{i}}\Big([0,1)^{2}).\end{cases}$
Let us fix $n\in\mathbb{N}$ and $(i_{1},...,i_{n})\in\left\{1,...,m\right\}^{n}.$ Notice that, for any $k\leq n$,
$f_{(i_{1},...,i_{n+k})}^{-1}(R_{\underline{i}})(x)=b^{k}(x)+\Big(-b^{-n(\tau_{1}-1)+k},b^{-n(\tau_{1}-1)+k}\Big)\times\Big(-b^{-n(\tau_{2}-1)+k},b^{-n(\tau_{2}-1)+k}\Big).$
Fix $\varepsilon,\varepsilon_{0}>0,$ By Proposition 4.9 applied simultaneously with $\tau_{1}=\tau_{1}-1,\tau_{2}=\tau_{2}-1$ and $\tau_{1}=\tau_{1}-1-\varepsilon_{0},$ $\tau_{2}=\tau_{2}-1-\varepsilon_{0}$, there exists $\rho>0$ such that, writing
$A_{\rho}=\left\{x\in K:\forall r\leq\rho,\begin{cases}r^{\varepsilon}\leq\frac{\mathcal{H}^{s}_{\infty}\Big((x+(-r^{\tau_{1}-1},r^{\tau_{1}-1})\times(-r^{\tau_{2}-1},r^{\tau_{2}-1}))\cap K\Big)}{r^{\max\left\{s(\tau_{1}-1),s(\tau_{2}-1)-(\tau_{2}-\tau_{1})(s_{0}-\alpha_{\nu})\right\}}}\\ r^{-\varepsilon}\geq\frac{\mathcal{H}^{s}_{\infty}\Big((x+(-r^{\tau_{1}-1-\varepsilon_{0}},r^{\tau_{1}-1-\varepsilon_{0}})\times(-r^{\tau_{2}-1-\varepsilon_{0}},r^{\tau_{2}-1-\varepsilon_{0}}))\cap K\Big)}{r^{\max\left\{s(\tau_{1}-1),s(\tau_{2}-1)-(\tau_{2}-\tau_{1})(s_{0}-\alpha_{\nu})\right\}}}\end{cases}\right\},$
one has $\nu(A_{\rho})\geq 1-\varepsilon_{0}.$ This implies, by ergodicity, that that for $\nu$-almost every $x$, there exists $k\in\mathbb{N}$ so large that for every $b^{k}x\in A_{\rho}.$
$b^{k}x+(-b^{-n(\tau_{1}-1)},b^{-n(\tau_{1}-1)})\times(-b^{-n(\tau_{2}-1)},b^{-n(\tau_{2}-1)})$
$\subset b^{k}x+\Big(-b^{-n(\tau_{1}-1)+k},b^{-n(\tau_{1}-1)+k}\Big)\times\Big(-b^{-n(\tau_{2}-1)+k},b^{-n(\tau_{2}-1)+k}\Big)$
and
$b^{k}x+(-b^{-n(\tau_{1}-1)+k},b^{-n(\tau_{1}-1)+k})\times(-b^{-n(\tau_{2}-1)},b^{-n(\tau_{2}-1)+k})$
$\subset b^{k}x+(-b^{-n(\tau_{1}-1-\varepsilon_{0})},b^{-n(\tau_{1}-1-\varepsilon_{0})})\times(-b^{-n(\tau_{2}-1-\varepsilon_{0})},b^{-n(\tau_{2}-1-\varepsilon_{0})}).$
Since
$\mathcal{H}^{s}_{\infty}(f_{(i_{1},...,i_{n+k})}^{-1}(R_{\underline{i}})\cap K)=b^{-(n+k)s}\times\mathcal{H}^{s}_{\infty}(R_{\underline{i}}\cap K)$
and, for every large enough $n$, one has
$b^{-n\varepsilon}\leq\frac{\mathcal{H}^{s}_{\infty}\Big(f_{(i_{1},...,i_{n+k})}^{-1}(R_{\underline{i}})\cap K\Big)}{b^{-n\max\left\{s(\tau_{1}-1),s(\tau_{2}-1)-(\tau_{2}-\tau_{1})(s_{0}-\alpha_{\nu})\right\}}}\leq b^{n\varepsilon},$
one gets that, provided that $n$ is large enough,
$b^{-2\varepsilon n}\leq\frac{\mathcal{H}^{s}_{\infty}(R_{\underline{i}}\cap K)}{b^{-ns}\times b^{-n\max\left\{s(\tau_{1}-1),s(\tau_{2}-1)-(\tau_{2}-\tau_{1})(s_{0}-\alpha_{\nu})\right\}}}\leq b^{2n\varepsilon}$
so that
$b^{-2\varepsilon n}\leq\frac{\mathcal{H}^{s}_{\infty}(R_{\underline{i}}\cap K)}{b^{-n\max\left\{\tau_{1},s\tau_{2}-(\tau_{2}-\tau_{1})(s_{0}-\alpha_{\nu})\right\}}}\leq b^{2n\varepsilon}.$
We are now ready to prove Theorem 3.4. Let
$\Lambda_{\mu}\subset\bigcup_{k\geq 1}\left\{1,...,m\right\}^{k}$
be a set of words such that there exists $z\in K$ for which
$\mu\Big(\limsup_{\underline{i}\in\Lambda_{\mu}}B\Big(f_{i}(z),|f_{\underline{i}}(K)|\Big)\Big)=1.$
Since for any $x\in K,$
$B\Big(f_{i}(z),|f_{\underline{i}}(K)|\Big)\Big)\subset B\Big(f_{i}(x),2|f_{\underline{i}}(K)|\Big)\Big),$
one has
$\mu\Big(\limsup_{\underline{i}\in\Lambda_{\mu}}B\Big(f_{i}(x),2|f_{\underline{i}}(K)|\Big)\Big)=1.$
In this, case (see [6]), for any $\varepsilon>0$ there exists a subset of words $\widetilde{\Lambda}_{\mu}\subset\Lambda_{\mu}$ such that
$\begin{cases}\forall\underline{i}\in\widetilde{\Lambda}_{\mu},\ \mu\Big(B\Big(f_{i}(x),2|f_{\underline{i}}(K)|\Big)\Big)\leq b^{-|\underline{i}|(\dim_{H}\mu-\varepsilon)}\\ \mu\Big(\limsup_{\underline{i}\in\widetilde{\Lambda}_{\mu}}B\Big(f_{i}(x),2|f_{\underline{i}}(K)|\Big)\Big)=1\end{cases}$
If $s_{\varepsilon}$ is solution to the equation
$\max\left\{s\tau_{1},s\tau_{2}-(\tau_{2}-\tau_{1})(s_{0}-\alpha_{\nu})\right\}+\varepsilon=\dim_{H}\mu-\varepsilon,$
for every $\underline{i}\in\widetilde{\Lambda}_{\mu}$ of large enough generation,
$\begin{cases}\mathcal{H}^{s_{\varepsilon}}_{\infty}(R_{\underline{i}}\cap K)\geq\mu\Big(B\Big(f_{i}(x),2|f_{\underline{i}}(K)|\Big)\Big)\\ \mu\Big(\limsup_{\underline{i}\in\widetilde{\Lambda}_{\mu}}B\Big(f_{i}(x),2|f_{\underline{i}}(K)|\Big)\Big)=1.\end{cases}$
Thus Theorem 4.5, one gets
$\dim_{H}\limsup_{\underline{i}\in\Lambda_{\mu}}R_{\underline{i}}\geq s_{\varepsilon}.$
Letting $\varepsilon\to 0,$ yields
$\dim_{H}\limsup_{\underline{i}\in\Lambda_{\mu}}R_{\underline{i}}\geq s$
where $s$ is solution to
$\max\left\{s\tau_{1},s\tau_{2}-(\tau_{2}-\tau_{1})(s_{0}-\alpha_{\nu})\right\}=\dim_{H}\mu,$
i.e.,
$s=\min\left\{\frac{\dim_{H}\mu}{\tau_{1}},\frac{\dim_{H}\mu+(\tau_{2}-\tau_{1})(s_{0}-\alpha_{\nu})}{\tau_{2}}\right\}.$
Assume in addition that
$\lim_{|\underline{i}|\to+\infty}\frac{-\log\mu\Big(f_{\underline{i}}([0,1]^{2}))\Big)}{|\underline{i}|\log b}=\dim_{H}\mu,$
Then, for any $\underline{i}\in\Lambda_{\mu},$ and any $\varepsilon>0,$ there exists $N\in\mathbb{N}$ such that for every $n\geq N,$ every $\underline{i}\in\Lambda_{\mu}\cap\left\{1,...,m\right\}^{n},$ writing $t_{\varepsilon}$ the solution to
$\max\left\{s\tau_{1},s\tau_{2}-(\tau_{2}-\tau_{1})(s_{0}-\alpha_{\nu})\right\}-\varepsilon=\dim_{H}\mu+\varepsilon,$
one has
$\mathcal{H}^{t_{\varepsilon}}_{\infty}(R_{\underline{i}}(x)\cap K)\leq b^{-n\varepsilon}\mu\Big(f_{\underline{i}}([0,1]^{d})\Big).$
And
$\sum_{n\geq 1}b^{-n\varepsilon}\sum_{\underline{i}\in\left\{1,...,m\right\}^{n}}\mu\Big(f_{\underline{i}}([0,1]^{d})\Big)=\sum_{n\geq 1}b^{-n\varepsilon}<+\infty.$
This yields in particular
$\sum_{\underline{i}\in\Lambda_{\mu}}\mathcal{H}^{t_{\varepsilon}}_{\infty}(R_{\underline{i}}(x)\cap K)<+\infty,$
so that $\dim_{H}\limsup_{\underline{i}\in\Lambda_{\mu}}R_{\underline{i}}(x)\leq t_{\varepsilon}.$
Letting $\varepsilon\to 0,$ one gets
$\dim_{H}\limsup_{\underline{i}\in\Lambda_{\mu}}R_{\underline{i}}\leq s,$
which finishes the proof.
5.2. Proof of Theorem 3.1
Let us fix $m,b\in\mathbb{N}$, $S=\left\{f_{1},...,f_{m}\right\}$ a two-dimensional base $b$ missing digit IFS, $\mu_{0}$ the corresponding Alfhors-regular measure and $s_{0}$ its dimension.
One will only prove Theorem 3.1 in the case where $(X_{n})_{n\in\mathbb{N}}$ is an i.i.d. sequence of law $\mu_{0}.$ In Section 5.2.3, we will explain how one recovers the result regards $\mu_{0}$-typical orbits by adapting a lemma established as[10, Lemma 7.3].
First, we justify, as in the previous section that, in the case where $\mu_{0}$ is carried by an horizontal or vertical line, Theorem 3.1 provides the same bound as in the case of random balls, thus is correct. In the case where $\mu_{0}$ is carried by an horizontal line, one has $\dim_{H}\pi_{2}\mu_{0}=0,$ so that one always as $\frac{1}{\tau_{1}}\leq s_{0}-\dim_{H}\pi_{2}\mu_{0}$ and $\dim_{H}W_{\tau_{1},\tau_{2}}=\frac{1}{\tau_{1}}.$
In the case where $\mu_{0}$ is carried by an horizontal line, $\dim_{H}\pi_{2}\mu_{0}=s_{0},$ so that $\kappa_{2}=s_{0}$ and $v_{\tau}(\kappa_{2})=0$, which yields
$\dim_{H}W_{\tau_{1},\tau_{2}}=\frac{1+(\tau_{2}-\tau_{1})\times 0}{\tau_{2}}=\frac{1}{\tau_{2}},$
and the conclusion of Theorem 3.1 holds true in that case.
We now focus on the case where $\mu_{0}$ is not carried bay an horizontal or vertical line and we use the same notations as in Section 4.3.
Given $\tau_{1},\tau_{2},r>0$ and $x=(x_{1},x_{2})\in\mathbb{R}^{2}$, let us write
$R_{\tau_{1},\tau_{2}}(x,r)=x+(-r^{\tau_{1}},r^{\tau_{1}})\times(-r^{\tau_{2}},r^{\tau_{2}}).$
Given $\alpha\in\mbox{Spectr}(\pi_{2}\mu_{0}),$ define $s_{\alpha}$ as the solution to
$\max\left\{s,s\tau-(\tau-1)(s_{0}-\alpha)\right\}=-(\tau-1)(\alpha_{0}-D_{\pi_{2}(\mu_{0})}(\alpha))+\frac{1}{\tau_{1}},$
i.e.
$s_{\alpha}=\min\left\{\frac{1-(\tau_{2}-\tau_{1})(\alpha-D_{\pi_{2}\mu_{0}}(\alpha))}{\tau_{1}},\frac{1+(\tau_{2}-\tau_{1})(s_{0}-2\alpha+D_{\pi_{2}\mu_{0}}(\alpha))}{\tau_{2}}\right\}.$
In the next sections, we show the following.
Theorem 5.1.
Let $(X_{n})_{n\in\mathbb{N}}$ be an i.i.d. sequence of common law $\mu_{0}$ and $\frac{1}{s_{0}}\leq\tau_{1}\leq\tau_{2}.$
Almost surely,
$\dim_{H}\limsup_{n\to+\infty}R_{\tau_{1},\tau_{2}}(X_{n},\frac{1}{n})=\max_{\alpha\in\mbox{Spectr}(\pi_{2}\mu_{0})}s_{\alpha}.$
We justify below that Theorem 5.1 implies Theorem 3.1.
Given $\alpha\in\mbox{ Spectr }(\pi_{2}\mu_{0}),$ notice that, writing again $\tau=\frac{\tau_{2}}{\tau_{1}},$ and assuming $\tau>1$,
$\frac{1-(\tau_{2}-\tau_{1})(\alpha-D_{\pi_{2}\mu_{0}}(\alpha))}{\tau_{1}}\leq\frac{1+(\tau_{2}-\tau_{1})(s_{0}-2\alpha+D_{\pi_{2}\mu_{0}}(\alpha))}{\tau_{2}}$
$\Leftrightarrow\frac{1}{\tau_{1}}-(\tau-1)(\alpha-D_{\pi_{2}\mu_{0}}(\alpha))\leq\frac{\frac{1}{\tau_{1}}+(\tau-1)(s_{0}-2\alpha+D_{\pi_{2}\mu_{0}}(\alpha))}{\tau}$
$\Leftrightarrow(\tau-1)\frac{1}{\tau_{1}}-\tau(\tau-1)(\alpha-D_{\pi_{2}\mu_{0}}(\alpha))\leq(\tau-1)(s_{0}-2\alpha+D_{\pi_{2}\mu_{0}}(\alpha))$
$\Leftrightarrow\frac{1}{\tau_{1}}\leq s_{0}+(\tau-2)\alpha-(\tau-1)D_{\pi_{2}\mu_{0}}(\alpha)$
$\Leftrightarrow\frac{1}{\tau_{1}}\leq v_{\tau}(\alpha).$
We consider $3$ cases separately.
Case 1: $\frac{1}{\tau}\leq v_{\tau}(\dim_{H}\pi_{2}(\mu_{0}))=s_{0}-\dim_{H}\pi_{2}\mu_{0}.$
Notice that $\alpha\mapsto v_{\tau}(\alpha)$ is convex, since $\alpha\mapsto D_{\pi_{2}\mu_{0}}(\alpha)$ is concave and its minimum is attained in $\alpha$ such that
$D^{\prime}_{\pi_{2}\mu_{0}}(\alpha)=\frac{\tau-2}{\tau-1}\leq 1,$
i.e. $\alpha=\kappa_{\frac{\tau-2}{\tau-1}}.$ As $\alpha\mapsto D^{\prime}_{\pi_{2}\mu_{0}}(\alpha)$ is non increasing and $D^{\prime}_{\pi_{2}\mu_{0}}(\dim_{H}\pi_{2}\mu_{0})=1,$ one has
$\kappa_{2}\leq\dim_{H}\pi_{2}\mu_{0}\leq\kappa_{\frac{\tau-2}{\tau-1}}.$
Moreover $\alpha\mapsto\frac{1-(\tau_{2}-\tau_{1})(\alpha-D_{\pi_{2}\mu_{0}}(\alpha))}{\tau_{1}}$ is non decreasing on $(-\infty,\dim_{H}\pi_{2}\mu_{0}],$ non increasing on $\mu_{0},+\infty)$ and reaches its maximum $M_{1}$ on $\dim_{H}\pi_{2}\mu_{0}$ with $M_{1}=\frac{1}{\tau_{1}}.$
Similarly, $\alpha\mapsto\frac{1+(\tau_{2}-\tau_{1})(s_{0}-2\alpha+D_{\pi_{2}\mu_{0}}(\alpha))}{\tau_{2}}$ is concave, non decreasing on $(-\infty,\kappa_{2}],$ non increasing on $\kappa_{2},+\infty)$ and reaches its maximum $M_{2}$ on $\kappa_{2}$ with
$M_{2}=\frac{1+(\tau_{2}-\tau_{1})(s_{0}-2\kappa_{2}+D_{\pi_{2}\mu_{0}}(\kappa_{2}))}{\tau_{2}}.$
This yields $v_{\tau}(\alpha)\geq\frac{1}{\tau_{1}}$ for any $\alpha\leq\dim_{H}\pi_{2}\mu_{0},$ so that
$\sup_{\alpha\in\mbox{ Spectr}(\pi_{2}\mu_{0})}s_{\alpha}=M_{1}=\frac{1}{\tau_{1}}.$
Case 2: $v_{\tau}(\dim_{H}\pi_{2}(\mu_{0}))\leq\frac{1}{\tau_{1}}\leq v_{\tau}(\kappa_{2}).$
Call $\beta_{\tau_{1},\tau_{2}}$ the solution to
$v_{\tau}(\beta_{\tau_{1},\tau_{2}})=\frac{1}{\tau_{1}}.$
The mapping $\alpha\mapsto\frac{1+(\tau_{2}-\tau_{1})(s_{0}-2\alpha+D_{\pi_{2}\mu_{0}}(\alpha))}{\tau_{2}}$ is non increasing on $\kappa_{2},\beta_{\tau_{1},\tau_{2}}]$ while $\alpha\mapsto\frac{1-(\tau_{2}-\tau_{1})(\alpha-D_{\pi_{2}\mu_{0}}(\alpha))}{\tau_{1}}$ is non decreasing on $\beta_{\tau_{1},\tau_{2}},+\infty).$ We conclude in that case that
$\sup_{\alpha\in\mbox{ Spectr}(\pi_{2}\mu_{0})}s_{\alpha}$ $=\frac{1-(\tau_{2}-\tau_{1})(\beta_{\tau_{1},\tau_{2}}-D_{\pi_{2}\mu_{0}}(\beta_{\tau_{1},\tau_{2}}))}{\tau_{1}}$
$=\frac{1+(\tau_{2}-\tau_{1})(s_{0}-2\beta_{\tau_{1},\tau_{2}}+D_{\pi_{2}\mu_{0}}(\beta_{\tau_{1},\tau_{2}}))}{\tau_{2}}$
Case 3: $\frac{1}{\tau}\geq v_{\tau}(\kappa_{2}).$
As both $\alpha\mapsto\frac{1+(\tau_{2}-\tau_{1})(s_{0}-2\alpha+D_{\pi_{2}\mu_{0}}(\alpha))}{\tau_{2}}$ and $\alpha\mapsto\frac{1+(\tau_{2}-\tau_{1})(s_{0}-2\alpha+D_{\pi_{2}\mu_{0}}(\alpha))}{\tau_{2}}$ are non decreasing on $(-\infty,\kappa_{2})$ and $\alpha\mapsto\frac{1+(\tau_{2}-\tau_{1})(s_{0}-2\alpha+D_{\pi_{2}\mu_{0}}(\alpha))}{\tau_{2}}$ is non increasing on $\kappa_{2},+\infty),$ $s_{\alpha}$ reaches its maximum on $\kappa_{2}$ so that
$\sup_{\alpha\in\mbox{ Spectr}(\pi_{2}\mu_{0})}s_{\alpha}$ $=\frac{1+(\tau_{2}-\tau_{1})(s_{0}-2\kappa_{2}+D_{\pi_{2}\mu_{0}}(\kappa_{2}))}{\tau_{2}}$
5.2.1. Proof of the upper-bound
Write
$R_{k}=R_{\tau_{1},\tau_{2}}(X_{k},\frac{1}{k})$
Our strategy to establish that, almost surely, one has
$\dim_{H}\limsup_{k\to+\infty}R_{k}\leq\sup_{\alpha\in\mbox{Spectr}(\pi_{2}\mu_{0})}s_{\alpha}$
is to show that, for every $\varepsilon>0,$ writing $s_{\varepsilon}=\sup_{\alpha\in\mbox{Spectr}(\pi_{2}\mu_{0})}s_{\alpha}+\varepsilon,$ one has
$\sum_{k\geq 1}\mathcal{H}^{s_{\varepsilon}}_{\infty}\Big(R_{k}\cap K\Big)<+\infty.$
Let us fix $\varepsilon>0.$ We recall that $\alpha\mapsto D_{\pi_{2}(\mu)}(\alpha)$ is continuous. Thus, by taking $\varepsilon$ smaller if one must, one may assume that $|\alpha-\alpha^{\prime}|\leq\varepsilon\Rightarrow|D_{\pi_{2}(\mu)}(\alpha)-D_{\pi_{2}(\mu)}(\alpha^{\prime})|\leq\varepsilon.$ Set
$\begin{cases}\alpha_{\min}=\inf\left\{\alpha:\ D_{\pi_{2}(\mu)}(\alpha)>0\right\}\\ \alpha_{\max}=\sup\left\{\alpha:\ D_{\pi_{2}(\mu)}(\alpha)>0\right\}.\end{cases}$
We recall that, since $\pi_{2}(\mu)$ is a non atomic self-similar measure, it is known that $0<\alpha_{\min}\leq\alpha_{\max}<+\infty.$ For $0\leq k\leq\lfloor\frac{\alpha_{\max}-\alpha_{\min}}{\varepsilon}\rfloor$, set
$I_{k}=[\alpha_{\min}+k\varepsilon,\alpha_{\min}+(k+1)\varepsilon[.$
We also write, for $k=0,...,\lfloor\frac{\alpha_{\max}-\alpha_{\min}}{\varepsilon}\rfloor,$ $\alpha_{k}=\alpha_{\min}+k\varepsilon$ and $\alpha_{\lfloor\frac{\alpha_{\max}-\alpha_{\min}}{\varepsilon}\rfloor+1}=\alpha_{\max}.$
Write also $\theta_{0}=-\frac{\log(p_{0})}{\log b}$ and $\theta_{1}=-\frac{\log(p_{b-1})}{\log b}.$ By changing slightly $\varepsilon$ if one must, we may assume that $\theta_{0},\theta_{1}\notin\left\{\alpha_{i}\right\}_{1\leq k\leq\lfloor\frac{\alpha_{\max}-\alpha_{\min}}{\varepsilon}\rfloor}.$
The following argument follows from Theorem 2.2 applied to $\pi_{2}(\mu_{0})$ (which satisfies the open set condition).
Lemma 5.2.
Let $\alpha_{\min}\leq\alpha\leq\alpha_{\max}$ be a real number and, given $r>0$ let us write
$\mathcal{P}_{\alpha}(r,\varepsilon)=\sup\#\left\{\mathcal{T}\right\},$
where $\mathcal{T}$ is a $r$-packing of $\mathrm{supp}(\pi_{2}(\mu_{0}))$ with, for every $B\in\mathcal{T}$,
$|B|^{\alpha+\varepsilon}\leq\pi_{2}\mu_{0}(B)\leq|B|^{\alpha-\varepsilon}.$
There exists $r_{\alpha}>0$ such that for every $r\leq r_{\alpha},$ one has
$r^{-D_{\pi_{2}(\mu_{0})}(\alpha)+\varepsilon}\leq\mathcal{P}_{\alpha}(r,\varepsilon)\leq r^{-D_{\pi_{2}(\mu_{0})}(\alpha)-\varepsilon}.$
Let us fix first $N$ large enough so that for every $0\leq i\leq\lfloor\frac{\alpha_{\max}-\alpha_{\min}}{\varepsilon}\rfloor+1,$ $r_{\alpha_{i}}\geq 2^{-N+1}.$ In addition, for $n\in\mathbb{N},$ we fix a packing $\mathcal{P}_{i,n,\varepsilon}$ of
$E_{i,n,\varepsilon}:=\left\{x:\ |B(x,2^{-n})|^{\alpha_{i}+\varepsilon}\leq\pi_{2}(\mu_{0})(B(x,2^{-n}))\leq|B(x,2^{-n})|^{\alpha_{i}-\varepsilon}\right\}$
by balls centered on $E_{i,n,\varepsilon}$ and of radius $2^{-n-2}.$ Notice that
$E_{i,n,\varepsilon}\subset\bigcup_{B\in\mathcal{P}_{i,N,\varepsilon}}2B$
and that
$\bigcup_{0\leq i\leq\lfloor\frac{\alpha_{\max}-\alpha_{\min}}{\varepsilon}\rfloor+1}E_{i,n,\varepsilon}=\pi_{2}(K).$
Let $D$ be a $b$-adic cude with $\mu_{0}(D)=|D|^{s_{0}}$ and $2^{\frac{n-1}{\tau_{1}}}\leq k\leq 2^{\frac{n}{\tau_{1}}}$ such that $X_{k}\in D.$
In order to estimates $\mathcal{H}^{s}_{\infty}\Big(R_{k}\cap K\Big)$, for $s\geq 0,$ we will distinguish two cases:
Case 1: $X_{k}$ is near the boundary of $D$
Recall that for $i=1,0,$
$\alpha_{i}=\lim_{r\to 0^{+}}\frac{\log\pi_{2}\mu_{0}\Big(B(i,r)\Big)}{\log r}=\frac{-\log p_{i}}{\log b}$
was assumed to satisfy $\alpha_{0}\leq\alpha_{1}.$
Let $X_{k}$ be such that
$X_{k}\in f_{D}\Big([0,1]\times[0,b^{-(\tau-1)n}]\cup[0,1]\times[1-b^{-(\tau-1)n},1]\Big).$
In this case, if $R_{k}:=R_{\tau_{1},\tau_{2}}(X_{k},\frac{1}{k})$ intersects $D$ and an other cube $D^{\prime}$ of generation $n$, one has
$R_{k}\subset S_{D,D^{\prime}}:=$ $f_{D}\Big([0,1]\times[0,b^{-(\tau-1)(n-1)}]\cup[0,1]\times[1-b^{-(\tau-1)(n-1)},1]\Big)$
$\bigcup f_{D^{\prime}}\Big([0,1]\times[0,b^{-(\tau-1)(n-1)}]\cup[0,1]\times[1-b^{-(\tau-1)(n-1)},1]\Big).$
so that
$\mathcal{H}^{s}_{\infty}(R_{k}\cap K)\leq b^{-(n-1)\max\left\{s,s\tau-(\tau-1)(s_{0}-\alpha_{0})\right\}}.$
And, since the coding of 0 has only 0’s as digits, by self-similarity,
$\mu_{0}\Big(S_{D,D^{\prime}}\Big)\leq 2\times\mu_{0}(D)\times b^{-(\tau-1)(n-1)\alpha_{0}},$
one has
$E[\#\left\{2^{\frac{n-1}{\tau_{1}}}\leq k\leq 2^{\frac{n}{\tau_{1}}}:\ X_{k}\in S_{D,D^{\prime}}\right\}]\leq 2^{\frac{n}{\tau_{1}}}\times 2\mu_{0}(D)\times b^{-(\tau-1)(n-1)\alpha_{0}}.$
This yields that
$E[\#\left\{2^{\frac{n-1}{\tau_{1}}}\leq k\leq 2^{\frac{n}{\tau_{1}}}:\ X_{k}\in\bigcup_{D,D^{\prime}}S_{D,D^{\prime}}\right\}]\leq C\times b^{-n((\tau-1)\alpha_{0}-\frac{1}{\tau_{1}})}.$
By Markov’s inequality,
$\mathbb{P}\Big(\#\left\{b^{\frac{n-1}{\tau_{1}}}\leq k\leq b^{\frac{n}{\tau_{1}}}:\ X_{k}\in\bigcup_{D,D^{\prime}}S_{D,D^{\prime}}\right\}\geq b^{n\varepsilon}b^{-n((\tau-1)\alpha_{0}-\frac{1}{\tau_{1}})}\Big)\leq b^{-n\varepsilon},$
which, by Borel cantelli lemma, implies that almost surely, there exists $n$ large enough so that
$\left\{b^{\frac{n-1}{\tau_{1}}}\leq k\leq b^{\frac{n}{\tau_{1}}}:\ X_{k}\in\bigcup_{D,D^{\prime}}S_{D,D^{\prime}}\right\}\leq b^{-n((\tau-1)\alpha_{0}-\frac{1}{\tau_{1}}-\varepsilon)}.$
Notice that if $(\tau-1)\alpha_{0}-\frac{1}{\tau_{1}}-\varepsilon>0,$ then almost surely, for $n$ large enough,
$\left\{b^{\frac{n-1}{\tau_{1}}}\leq k\leq b^{\frac{n}{\tau_{1}}}:\ X_{k}\in\bigcup_{D,D^{\prime}}S_{D,D^{\prime}}\right\}=0,$
so that we may assume that $(\tau-1)\alpha_{0}-\frac{1}{\tau_{1}}-\varepsilon\leq 0,$
Thus if $s$ is solution to $\max\left\{s,s\tau-(\tau-1)(s_{0}-\alpha_{0})\right\}=-(\tau-1)\alpha_{0}+\frac{1}{\tau_{1}}+\varepsilon,$ one has for every $\kappa>1,$
$\sum_{X_{k}\in\bigcup_{D,D^{\prime}}S_{D,D^{\prime}}}\mathcal{H}^{s\kappa}_{\infty}\Big(R_{k}\cap K\Big)\leq\sum_{n\geq 1}b^{-n((\tau-1)\alpha_{0}-\frac{1}{\tau}-\varepsilon)}b^{n\kappa((\tau-1)\alpha_{0}-\frac{1}{\tau}-\varepsilon))}<+\infty.$
We conclude that
$\dim_{H}\limsup_{X_{k}\in\bigcup_{D,D^{\prime}}S_{D,D^{\prime}}}R_{k}\leq\kappa s$
for any $\kappa,$ so that
$\dim_{H}\limsup_{X_{k}\in\bigcup_{D,D^{\prime}}S_{D,D^{\prime}}}R_{k}\leq s.$
In addition, recall that for any $\alpha\in\mbox{Spectr}(\pi_{2}\mu_{0})$, $\alpha-D_{\pi_{2}\mu_{0}}\geq 0$ and that $s_{\alpha_{0}}$ is solution to the equation
$\max\left\{s,s\tau-(\tau-1)(s_{0}-\alpha_{0})\right\}=-(\tau-1)(\alpha_{0}-D_{\pi_{2}(\mu_{0})}(\alpha_{0}))+\frac{1}{\tau_{1}}.$
This yields $s\leq s_{\alpha_{0}}+\varepsilon$ provided that $\varepsilon$ was chosen small enough to begin with.
Case 2: $X_{k}$ is not near the boundary of $D$
Assume now that $X_{k}\in D\setminus f_{D}\Big([0,1]\times[0,b^{(\tau-1)n}]\cup[0,1]\times[1-b^{(\tau-1)n},1]\Big)$ and recall that, by self-similarity, one has
$\mathcal{H}^{s}_{\infty}(R\cap K\cap D)\leq\mathcal{H}^{s}_{\infty}(R\cap K)\leq 3\mathcal{H}^{s}_{\infty}(R\cap K\cap D).$
Recall that
$D\cap K\subset\bigcup_{0\leq i\leq\frac{\alpha_{\max}-\alpha_{\min}}{\varepsilon}\rfloor+1}f_{D}\circ\pi_{2}^{-1}\Big(\bigcup_{B\in\mathcal{P}_{i,\lfloor(\tau-1)n\rfloor-1,\varepsilon}}B\Big).$
Fix $0\leq i\leq\lfloor\frac{\alpha_{\max}-\alpha_{\min}}{\varepsilon}\rfloor+1$ and assume that $X_{k}\in f_{D}\circ\pi_{2}^{-1}(B),$ where $B\in\mathcal{P}_{i,\lfloor(\tau-1)n\rfloor-1,\varepsilon}.$ By definition of $\mathcal{P}_{i,\lfloor(\tau-1)n\rfloor-1,\varepsilon}$ and by self-similarity, one has
$\mu_{0}(f_{D}\circ\pi_{2}^{-1}(2B))\leq(2|B|)^{\alpha_{i}-\varepsilon}\times\mu_{0}(D).$
In addition, if $C$ is a $b$-adic cube of generation $\lfloor n\tau\rfloor$ intersecting $R\cap D\cap K,$ then $C\subset f_{D}\circ\pi_{2}^{-1}(2B).$ We conclude that
$\#\left\{C\in\mathcal{D}_{\lfloor n\tau\rfloor}:C\cap R\cap D\cap K\neq\emptyset\right\}$ $\leq\frac{\mu_{0}\Big(f_{D}\circ\pi_{2}^{-1}(2B)\Big)}{b^{-\lfloor n\tau\rfloor s_{0}}}\leq Cb^{n\tau s_{0}-ns_{0}-(\tau-1)(\alpha_{i}-\varepsilon)}$
$\leq Cb^{n(\tau-1)(s_{0}-\alpha_{i}+\varepsilon)}.$
By covering $R\cap D\cap K$ by $b$-adic cubes of generation $\lfloor n\tau\rfloor$ and $n$, we obtain that
$\mathcal{H}^{s}_{\infty}(R\cap D\cap K)\leq b^{-n\max\left\{s,s\tau-(\tau-1)(s_{0}-\alpha_{i}+\varepsilon)\right\}}.$
Moreover, by self-similarity,
$\mu_{0}\Big(f_{D}\circ\pi_{2}^{-1}\Big(\bigcup_{B\in\mathcal{P}_{i,\lfloor(\tau-1)n\rfloor-1,\varepsilon}}B\Big)\Big)$
$\leq\mu_{0}(D)\times\#\left\{\mathcal{P}_{i,\lfloor(\tau-1)n\rfloor-1,\varepsilon}\right\}\times\Big(2|B|\Big)^{\alpha_{i}-\varepsilon}\leq C\mu_{0}(D)\times b^{-n((\tau-1)(\alpha_{i}-\varepsilon)-(\tau-1)(D_{\pi_{2}(\mu_{0})}+\varepsilon))}$
$\leq C\mu_{0}(D)b^{-(\tau-1)n(\alpha_{i}-D_{\pi_{2}(\mu_{0})}(\alpha_{i})-2\varepsilon)}.$
Writing
$\begin{cases}G_{n,i,\varepsilon}=\bigcup_{D\in\mathcal{D}_{n}}f_{D}\circ\pi_{2}^{-1}\Big(\bigcup_{B\in\mathcal{P}_{i,\lfloor(\tau-1)n\rfloor-1,\varepsilon}}B\Big)\setminus f_{D}\Big([0,1]\times[0,b^{(\tau-1)n}]\cup[0,1]\times[1-b^{(\tau-1)n},1]\Big)\\ U_{n,i,\varepsilon}=\left\{b^{\frac{(n-1)}{\tau_{1}}}\leq k\leq b^{\frac{n}{\tau_{1}}}:\ X_{k}\in G_{n,i,\varepsilon}\right\}\end{cases}$
$E[\#\left\{U_{n,i,\varepsilon}\right\}]\leq b^{\frac{n}{\tau_{1}}}\times Cb^{-(\tau-1)n(\alpha_{i}-D_{\pi_{2}(\mu_{0})}(\alpha_{i})-2\varepsilon)}=Cb^{-n(-\frac{1}{\tau_{1}}+(\tau-1)(\alpha_{i}-D_{\pi_{2}(\mu_{0})}(\alpha_{i})-2\varepsilon))}.$
Thus, by Markov’s inequality and Borel-Cantelli lemma, almost surely, there exists $n_{i,\varepsilon}$ large enough so that
$\left\{U_{n,i,\varepsilon}\right\}\leq Cb^{-n(-\frac{1}{\tau_{1}}+(\tau-1)(\alpha_{i}-D_{\pi_{2}(\mu_{0})}(\alpha_{i})-3\varepsilon))}.$
Using the same argument as in case $(1)$, one assume that $\alpha_{i}$ is such that
$\frac{1}{\tau_{1}}+(\tau-1)(\alpha_{i}-D_{\pi_{2}(\mu_{0})}(\alpha_{i})-3\varepsilon)\leq 0.$
Let $s$ be the solution to
$\max\left\{s,s\tau-(\tau-1)(s_{0}-\alpha_{i}+\varepsilon)\right\}=\frac{1}{\tau_{1}}-(\tau-1)(\alpha_{i}-D_{\pi_{2}(\mu_{0})}(\alpha_{i})-3\varepsilon).$
The same argument in case $(1)$ yields that for every $\kappa>1$
$\sum_{n\geq 1,k\in U_{n,i,\varepsilon}}\mathcal{H}^{\kappa s}_{\infty}(R_{k}\cap K)<+\infty,$
so that
$\dim_{H}\limsup_{k\in U_{n,i,\varepsilon}}R_{k}\leq s.$
In addition, since $s_{\alpha_{i}}$ is solution to
$\max\left\{s,s\tau-(\tau-1)(s_{0}-\alpha_{i})\right\}=\frac{1}{\tau_{1}}-(\tau-1)(\alpha_{i}-D_{\pi_{2}(\mu_{0})}(\alpha_{i})),$
one has $s\leq s_{\alpha_{i}}+\varepsilon\leq\sup_{\alpha}s_{\alpha}+\varepsilon$ provided that $\varepsilon$ was chosen small enough to begin with.
5.2.2. Proof of the lower-bound
Recall that, for $\alpha\in\mbox{Spectr}(\pi_{2}\mu_{0}),$
$s_{\alpha}=\min\left\{\frac{1-(\tau_{2}-\tau_{1})(\alpha-D_{\pi_{2}\mu_{0}}(\alpha))}{\tau_{1}},\frac{1+(\tau_{2}-\tau_{1})(s_{0}-2\alpha+D_{\pi_{2}\mu_{0}}(\alpha))}{\tau_{2}}\right\}.$
The following fact is a consequence of Proposition 2.1
Proposition 5.3.
There exists $\nu_{\alpha}$, $\times b$-ergodic (hence exact-dimensional) such that
  • (1)
    $\nu_{\alpha}\Big(\left\{x:\lim_{r\to 0^{+}}\frac{-\log\pi_{2}\mu_{0}(D_{n}(x))}{n\log b}=\alpha\right\}\Big)=1,$
  • (2)
    $\dim_{H}\nu_{\alpha}=D_{\pi_{2}\mu_{0}}(\alpha).$
By Proposition 5.3, there exists $N\in\mathbb{N}$ such that the set
$F_{N,\varepsilon,\alpha}=\left\{x:\ \forall n\geq N,\begin{cases}b^{-n(D_{\pi_{2}\mu_{0}}(\alpha)+\varepsilon)}\leq\nu_{\alpha}(D_{n}(x))\leq b^{-n(D_{\pi_{2}\mu_{0}}(\alpha)-\varepsilon)}\\ b^{-n(\alpha+\varepsilon)}\leq\pi_{2}\mu_{0}(D_{n}(x))\leq b^{-n(\alpha-\varepsilon)}\end{cases}\right\}$
satisfies
$\nu_{\alpha}(F_{N,\varepsilon,\alpha})\geq\frac{1}{2}.$
Notice that, for any $n\in\mathbb{N}$ and $x\in\mathrm{supp}(\nu_{\alpha})\subset\left\{0\right\}\times[0,1],$
$\pi_{2}^{-1}(D_{n}(x))\cap K=[0,1]\times I\cap K,$
where $\left\{0\right\}\times I=D_{n}(x)\cap\left\{0\right\}\times[0,1].$ Hence,
$C^{-1}\leq\frac{\mathcal{H}^{s}_{\infty}([0,1]\times I\cap K)}{\inf_{0\leq\gamma\leq 1}b^{\gamma ns_{0}}\times\pi_{2}\mu_{0}(D_{\lfloor\gamma n\rfloor}(x))b^{-n\gamma s}}\leq C.$
Thus, for any $x\in F_{N,\alpha,\varepsilon},$ for any $n\geq N$ and $\gamma\geq\frac{n}{N},$ one has
$b^{-\gamma n(\alpha+\varepsilon)}\leq\pi_{2}\mu_{0}(D_{\lfloor\gamma n\rfloor}(x))\leq b^{-\gamma n(\alpha-\varepsilon)}$
and there exists $\kappa_{N}>0$ such that
$\kappa_{N}^{-1}\leq\inf_{\frac{n}{N}\leq\gamma\leq 1}b^{-\gamma ns_{0}}\times\pi_{2}\mu_{0}(D_{\lfloor\gamma n\rfloor}(x))b^{-n\gamma s}\leq\kappa_{N}.$
This yields
$\inf_{0\leq\gamma\leq\frac{n}{N}}b^{-\gamma n(s-s_{0}+\alpha+\varepsilon)}\leq\inf_{0\leq\gamma\leq\frac{n}{N}}b^{\gamma ns_{0}}\times\pi_{2}\mu_{0}(D_{\lfloor\gamma n\rfloor}(x))b^{-n\gamma s}\leq\inf_{0\leq\gamma\leq\frac{n}{N}}b^{-\gamma n(s-s_{0}+\alpha-\varepsilon)}$
If $s\geq s_{0}-\alpha+\varepsilon$ or $s\leq s_{0}-\alpha-\varepsilon,$ both $\gamma\mapsto\gamma(s-(s_{0}-\alpha-\varepsilon))$ and $\gamma\mapsto\gamma(s-(s_{0}-\alpha+\varepsilon))$ are non increasing or non decreasing. We conclude that, for any such $s$,
$\frac{1}{C\log n}\min\left\{\kappa_{N},b^{-n(s-s_{0}+\alpha+\varepsilon)}\right\}\leq\mathcal{H}^{s}_{\infty}([0,1]\times I\cap K)\leq C\log(n)\min\left\{\kappa_{N},b^{-n(s-s_{0}+\alpha-\varepsilon)}\right\}.$
Moreover, since the content is non increasing in $s$, for any $s_{0}-\alpha-\varepsilon\leq s\leq s_{0}-\alpha+\varepsilon,$ one has
$\frac{1}{C\log n}\min\left\{\kappa_{N},b^{-n(s-s_{0}+\alpha+2\varepsilon)}\right\}\leq\mathcal{H}^{s}_{\infty}([0,1]\times I\cap K)\leq C\log(n)\min\left\{\kappa_{N},b^{-n(s-s_{0}+\alpha-2\varepsilon)}\right\}$
so that the latest inequalities holds for any $s\leq s_{0}.$ From now on, we consider $n$ large enough so that $(\tau-1)n\geq N+1$ and
$\mathcal{P}_{(\tau-1)n}=\left\{D_{\lfloor(\tau-1)n\rfloor+1}(x),x\in F_{N,\varepsilon,\alpha}\right\}.$
It follows from the definition of $F_{N,\varepsilon,\alpha}$ that
$\begin{cases}\#\mathcal{P}_{(\tau-1)n}\geq b^{(\tau-1)n(D_{\pi_{2}\mu_{0}}(\alpha)-\varepsilon)}\\ \forall I\in\mathcal{P}_{(\tau-1)n},\ b^{-n(\tau-1)(\alpha+\varepsilon)}\leq\pi_{2}\mu_{0}(D_{n}(x))\leq b^{-n(\tau-1)(\alpha-\varepsilon)}.\end{cases}$
Hence, we obtain
$\pi_{2}\mu_{0}\Big(\bigcup_{I\in\mathcal{P}_{(\tau-1)n}}I\Big)\geq b^{-n(\tau-1)(\alpha-D_{\pi_{2}\mu_{0}}(\alpha)+2\varepsilon)}.$
Let us write
$\delta_{\alpha}=\frac{s_{0}}{\frac{1}{\tau_{1}}-(\tau-1)(\alpha-D_{\pi_{2}\mu_{0}}(\alpha))}.$
Notice that, since $\tau_{1}\geq\frac{1}{s_{0}}$, $\tau\geq 1$ and $\alpha-D_{\pi_{2}\mu_{0}}(\alpha)\geq 0,$ one has $\delta_{\alpha}\geq 1.$
Fix $x\in K$ and set $k=\lfloor\frac{n}{\delta_{\alpha}+\varepsilon}\rfloor$ For any $D\subset D_{k}(x)$ with $D\in\mathcal{D}_{n},$ by self-similarity, one has
$\mu_{0}\Big(f_{D}\Big(\pi_{2}^{-1}\Big(\bigcup_{I\in\mathcal{P}_{(\tau-1)n}}I\Big)\Big)\Big)=\mu_{0}(D)\times\pi_{2}\mu_{0}\Big(\bigcup_{I\in\mathcal{P}_{(\tau-1)n}}I\Big).$
This yields that
$\mu_{0}\Big(\bigcup_{D\in\mathcal{D}_{n},D\subset D_{k}(x)}(f_{D}\Big(\pi_{2}^{-1}\Big(\bigcup_{I\in\mathcal{P}_{(\tau-1)n}}I\Big)\Big)\Big)=\mu(D_{k}(x))\times\pi_{2}\mu_{0}\Big(\bigcup_{I\in\mathcal{P}_{(\tau-1)n}}I\Big)$
$\geq b^{-n\Big(\frac{s_{0}}{(1+\varepsilon)\delta_{\alpha}}+(\tau-1)(\alpha-D_{\pi_{2}\mu_{0}}(\alpha)+2\varepsilon)\Big)}=b^{-n(\frac{1}{(1+\varepsilon)\tau_{1}})+(\tau-1)((1-\frac{1}{1+\varepsilon})(\alpha-D_{\pi_{2}\mu_{0}}(\alpha))+2\varepsilon)}\geq b^{\frac{-n}{(1+\frac{\varepsilon}{2})\tau_{1}}}$
provided that $\varepsilon$ was chosen small enough to begin with. Write
$A_{k}=\bigcup_{D\in\mathcal{D}_{n},D\subset D_{k}(x)}(f_{D}\Big(\pi_{2}^{-1}\Big(\bigcup_{I\in\mathcal{P}_{(\tau-1)n}}I\Big)\Big).$
Then, for any $p\in\mathbb{N}$ and any constant $C>0$ and $q\geq Cb^{\frac{n}{\tau_{1}}},$
(13) $\mathbb{P}\Big(X_{p},...,X_{p+q}\notin A_{k}\Big)\leq(1-\mu_{0}(A_{k}))^{Cb^{\frac{n}{\tau_{1}}}}\leq e^{-Cb^{\frac{n}{\tau_{1}}}\mu_{0}(A_{k})}.$
Since $b^{\frac{n}{\tau_{1}}}\mu_{0}(A_{k})\geq b^{\frac{n}{\tau_{1}}(1-\frac{1}{1+\varepsilon})}\to+\infty.$ Since $\left\{b^{\frac{n}{\tau_{1}}}<p\leq b^{\frac{n+1}{\tau_{1}}}\right\}\geq Cb^{\frac{n}{\tau_{1}}},$ by Borel-Cantelli we obtain that almost surely, for every large enough $k$, there exists $b^{\frac{n}{\tau_{1}}}\leq p\leq b^{\frac{n+1}{\tau_{1}}}$ such that $X_{p}\in A_{k}.$ Now, fixing $D\in\mathcal{D}_{n}$ such that $D\subset D_{k}(x),$ $I\in\mathcal{P}_{(\tau-1)n}$ and $X_{p}\in f_{D}(\pi_{2}^{-1}(I)).$ Notice that $f_{D}\Big(\pi_{2}^{-1}(I)\cap[0,1]^{2}\Big)$ is a rectangle, product of a $b$-adic interval of generation $n$ with a $b$-adic interval of generation $\tau n.$ Since $p\leq 2^{\frac{n-1}{\tau_{1}}},$ one has
$\begin{cases}\frac{1}{k^{\tau_{1}}}\geq 2^{-n}\\ \frac{1}{k^{\tau_{2}}}\geq 2^{-n\tau}\end{cases}$
which yields that $f_{D}\Big(\pi_{2}^{-1}(I)\cap[0,1]^{2}\Big)\subset R_{k}$ and, provided that $n$ is large enough,
$\mathcal{H}^{s}_{\infty}\Big(R_{k}\cap K\Big)$ $\geq\mathcal{H}^{s}_{\infty}\Big(f_{D}\Big(\pi_{2}^{-1}(I)\cap[0,1]^{2}\Big)\Big)$
$\geq\frac{C}{\log((\tau-1)n)}|D|^{s}\min\left\{\kappa_{N},b^{-n(\tau-1)(s-s_{0}+2\varepsilon)}\right\}$
$\geq\theta_{N}\min\left\{b^{-n(1+\varepsilon)s},b^{-n(\tau-1)(s-s_{0}+3\varepsilon)}\right\}$
$=\theta_{N}b^{-n\max\left\{s(1+\varepsilon),\tau s-(\tau-1)(s_{0}-\alpha-3\varepsilon)\right\}}.$
Also,
$s_{0}k=\frac{s_{0}n}{(1+\varepsilon)\delta_{\alpha}}=\frac{1}{1+\varepsilon}\times\Big(\frac{1}{\tau_{1}}-(\tau-1)(\alpha-D_{\pi_{2}(\mu_{0})}(\alpha))\Big)$
so that, if $s$ is solution to
(14) $\frac{1}{1+\varepsilon}\times\Big(\frac{1}{\tau_{1}}-(\tau-1)(\alpha-D_{\pi_{2}(\mu_{0})}(\alpha))\Big)=\max\left\{s(1+\varepsilon),\tau s-(\tau-1)(s_{0}-\alpha-3\varepsilon)\right\},$
for every $x$, almost surely, for every large enough $k\in\mathbb{N}$, there exists $p$ such that $X_{p}\in D_{k}(x)$ and
$\mathcal{H}^{s}_{\infty}(R_{p}\cap K)\geq\mu_{0}(D_{k}(x).$
Note also that
$D_{k}(x)\cup R_{p}\subset B\Big(X_{p},\frac{1}{p^{\frac{\tau_{1}}{(1+\varepsilon)\delta_{\alpha}}}}\Big)\text{ and }\frac{1}{p^{\frac{\tau_{1}}{(1+\varepsilon)\delta_{\alpha}}}}\leq b^{2}|D_{n}(x)|$
which implies in particular that $x\in B\Big(X_{p},\frac{1}{p^{\frac{\tau_{1}}{(1+\varepsilon)\delta_{\alpha}}}}\Big)$ and
$\mathcal{H}^{s}_{\infty}(R_{p}\cap K)\geq C\mu_{0}\Big(B\Big(X_{p},\frac{1}{p^{\frac{\tau_{1}}{(1+\varepsilon)\delta_{\alpha}}}}\Big)\Big)$
so that, almost surely,
$x\in\limsup_{p:\ \mathcal{H}^{s}_{\infty}(R_{p}\cap K)\geq C\mu_{0}\Big(B\Big(X_{p},\frac{1}{p^{\frac{\tau_{1}}{(1+\varepsilon)\delta_{\alpha}}}}\Big)\Big)}B\Big(X_{p},\frac{1}{p^{\frac{\tau_{1}}{(1+\varepsilon)\delta_{\alpha}}}}\Big).$
Thus, by Fubini,
$\int\mathbb{P}\Big(\limsup_{p:\ \mathcal{H}^{s}_{\infty}(R_{p}\cap K)\geq C\mu_{0}\Big(B\Big(X_{p},\frac{1}{p^{\frac{\tau_{1}}{(1+\varepsilon)\delta_{\alpha}}}}\Big)\Big)}B\Big(X_{p},\frac{1}{p^{\frac{\tau_{1}}{(1+\varepsilon)\delta_{\alpha}}}}\Big)\Big)d\mu_{0}(x)=1$
$\Leftrightarrow\int\mu_{0}\Big(\limsup_{p:\ \mathcal{H}^{s}_{\infty}(R_{p}\cap K)\geq C\mu_{0}\Big(B\Big(X_{p},\frac{1}{p^{\frac{\tau_{1}}{(1+\varepsilon)\delta_{\alpha}}}}\Big)\Big)}B\Big(X_{p},\frac{1}{p^{\frac{\tau_{1}}{(1+\varepsilon)\delta_{\alpha}}}}\Big)\Big)d\mathbb{P}=1$
and almost surely,
$\mu_{0}\Big(\limsup_{p:\ \mathcal{H}^{s}_{\infty}(R_{p}\cap K)\geq C\mu_{0}\Big(B\Big(X_{p},\frac{1}{p^{\frac{\tau_{1}}{(1+\varepsilon)\delta_{\alpha}}}}\Big)\Big)}B\Big(X_{p},\frac{1}{p^{\frac{\tau_{1}}{(1+\varepsilon)\delta_{\alpha}}}}\Big)\Big)=1.$
Theorem 4.5 yields
$\dim_{H}\limsup_{k\to+\infty}R_{k}\geq s.$
Recalling that $s$ satisfies (14) and that $s_{\alpha}$ is solution to the equation
$\Big(\frac{1}{\tau_{1}}-(\tau-1)(\alpha-D_{\pi_{2}(\mu_{0})}(\alpha))\Big)=\max\left\{s,\tau s-(\tau-1)(s_{0}-\alpha)\right\},$
one has $s\geq s_{\alpha}-\varepsilon$ provided that $\varepsilon$ was chosen small enough. Since, $\varepsilon,\alpha$ are arbitrary, Theorem 3.1 is proved.
5.2.3. Dealing with the case of orbits
In this section we explain how to adapt the proof made in the case where $(X_{n})_{n\in\mathbb{N}}$ is i.i.d. of law $\mu_{0}$ to the case where $X_{n}=T_{b}^{n}(x),$ where $x$ is $\mu_{0}$-typical and $T_{b}:\mathbb{T}^{2}\to\mathbb{T}^{2}$ is defined by $T_{b}(x)=bx$.
First, notice that the independency of the sequence $(X_{n})_{n\in\mathbb{N}}$ was not used to establish the upper-bound. Thus, the argument readily applies to $(Y_{n}=b^{n}x)_{n\in\mathbb{N}}.$
However, the independency is used while estimating the lower-bound in (13). With the same notations as in the corresponding subsection, the idenpendency was used to establish that given any $b$-adic cube $D\in\mathcal{D}_{\lfloor\frac{n}{(1+\varepsilon)\delta_{\alpha}}\rfloor}$ with $D\cap K\neq\emptyset$ if
$\mathcal{R}_{D}=\left\{D^{\prime}\in\mathcal{D}_{n},D^{\prime}\cap K\neq\emptyset f_{D^{\prime}}\Big(\pi_{2}^{-1}(B)\Big),B\in\mathcal{P}_{\lfloor(\tau-1)n\rfloor}\right\},$
then almost surely there exists $b^{\frac{n-1}{\tau_{1}}}\leq k\leq b^{\frac{n}{\tau_{1}}}$ such that $X_{k}\in\bigcup_{R\in\mathcal{R}_{D}}R.$
It is established in [10, Proposition 3.10] that there exists $C>0$ and $\tau_{1}$ such that for every ball AA and Borel set $B$, for every $n\in\mathbb{N},$ one has
$\mu_{0}\Big(T_{b}^{-n}(B)\cap A\Big)\leq\mu(A)\times\mu(B)+C\tau^{n}\mu(B).$
Thus, as $\left\{R_{D}\right\}$ is polynomial in $|D|,$ the same argument as in [10, Lemma 7.6] holds to show that the same conclusion as in the case where $(X_{n})_{n\in\mathbb{N}}$ was an i.i.d. sequence holds true.
6. Acknowledgments
The author would like to thank D.J. Feng and S.Seuret for sharing the random covering problem of self-similar sets by rectangles with him.